### 11 Mar 2026 v1

# Unified Flavor:

Abstract
[…]
Froggatt–Nielsen mechanism, vectorlike quarks, flavor hierarchy, discrete gauge symmetry, axion quality, CKM mixing, PMNS mixing, Yukawa textures, neutrino mass
I Introduction
The core of this Froggatt–Nielsen (FN) flavor phenomenology is quadrilateral combinations of fermion masses evaluated at scale $M_{Z}$ that are found to be quantized as power laws $B^{n}$ on a lattice with a universal constant $B=75/14$ and integer quantum numbers $n$. From these integers $n$, the masses and the weak mixing elements are inferred to be power laws $(1/B)^{p}$ where $\epsilon=1/B$ is the small FN parameter and the $p$-values are individually predicted as rational fractions that are determined by the $n$-values.
Understanding the origin of fermion mass hierarchies and flavor mixing remains a central challenge in particle physics. The Froggatt–Nielsen (FN) mechanism [1, 2, 3] generates hierarchical Yukawa couplings through powers of a small expansion parameter $\epsilon=\langle\Phi\rangle/\Lambda$ set by the vacuum expectation value (VEV) of a single flavon field $\Phi$. The physical picture is as follows. Each Standard Model fermion carries a generation-dependent $\mathrm{U}(1)_{\rm FN}$ charge (or, more generally, a discrete $Z_{N}$ charge). A Yukawa coupling $\bar{f}_{L}\,f_{R}\,H$ is forbidden at tree level by the charge mismatch between the left- and right-handed fermions; it is generated only through higher-dimension operators containing $n$ powers of $\Phi/\Lambda$, where $n$ is fixed by charge selection rules. After the flavon develops its VEV the effective Yukawa coupling scales as $y_{ij}\sim\epsilon^{n_{ij}}$, so that the observed mass hierarchies—spanning five orders of magnitude in the quark sector—are converted into modest differences in the integer (or, once discrete gauge symmetries are admitted, rational) exponents $n_{ij}$. In the standard FN framework these exponents are integers counting flavon insertions along a single messenger line. The key innovation of the $B$-lattice program is that discrete gauge symmetries of the type analyzed by Ibáñez and Ross [4] permit rational exponents quantized in ninths, enabling a single expansion parameter $\epsilon=1/B$ to fit both quark and lepton mass spectra with no additional free parameters beyond $\mathcal{O}(1)$ coefficients.
It is important to stress that the fractional powers appearing in the effective field theory (EFT) do not appear in the ultraviolet Lagrangian itself. The UV completion is fully renormalizable: one introduces heavy vectorlike messenger fermions with mass $M$, and the light Standard Model fields couple to these messengers and to the flavon $\Phi$ through ordinary Yukawa interactions [1, 2, 3]. Integrating out a chain of $n$ messenger fields generates an effective operator proportional to $(\langle\Phi\rangle/M)^{n}$ [5]. In the $B$-lattice framework the discrete $Z_{N}$ charge differences between adjacent chain sites need not be unity; they can take values such as $1$, 22, or $4$ (mod 9), so that each hop contributes a fractional exponent $h_{k}/9$ while still requiring an integer number of flavon insertions $h_{k}$ at that vertex. The “ninths” exponents are therefore a compact EFT bookkeeping for short messenger chains—typically three links—whose individual hops are governed by discrete gauge charge selection rules [4]. Sections III and IV below make this correspondence explicit.
It is worth emphasizing that the numerator of a ninths exponent is not the chain length: it is the sum of discrete charge differences along the chain. If one simply assigned the flavon a charge of $1/9$, then integrating out $n$ messengers would give $\epsilon^{n/9}$ with $n$ equal to the number of links. That would require 18 messengers to produce $|V_{cb}|\sim\epsilon^{2}$, or 27 for $|V_{ub}|\sim\epsilon^{3}$—hardly economical. In the $B$-lattice framework the flavon carries charge 1 under $Z_{9}$, but the messenger sites carry different $Z_{9}$ charges, so the charge gap $h_{k}$ at each nearest-neighbor hop need not be unity. For instance, a three-site chain with hop charges $(h_{1},h_{2},h_{3})=(1,2,4)\;\mathrm{mod}\;9$ yields a total exponent $(1+2+4)/9=7/9$ from only three links. Because the set $\{1,2,4\}$ generates all residues mod 9, different Yukawa entries—corresponding to different endpoint dressings on the same short chain—access the full ninths lattice $\{n/9\}$ without requiring long chains. The denominator 9 thus reflects the order of the discrete gauge group, while the numerator encodes the charge assignment pattern at the chain sites.
A series of recent papers [6, 5, 7, 8] has demonstrated that a specific quantization of the FN exponents—the $B$-lattice, with $B=75/14$ and $\epsilon=1/B\simeq 0.187$—organizes quark and lepton masses, CKM and Pontecorvo–Maki–Nakagawa–Sakata (PMNS) mixing, and CP violation into a unified framework.
Independently, theory-space locality and clockwork/chain constructions have shown that exponential hierarchies can emerge from nearest-neighbor couplings among vectorlike fermions [9, 10, 11]. In a recent development, Arkani-Hamed, Figueiredo, Hall, and Manzari (AFHM) [12] have proposed a concrete class of TeV-scale chain models in which locality in the VLQ chains simultaneously generates hierarchical Yukawa matrices and suppresses flavor- and CP-violating processes, with a natural extension to neutrino masses. Their framework shares the essential mechanism exploited here—nearest-neighbor VLQ chains producing effective Yukawa couplings as products of sequential mass mixings—but does not incorporate the discrete $\mathbb{Z}_{18}$ lattice quantization that underpins the $B$-lattice program and its connection to the axion quality problem. VLQs have also been studied extensively both as mediators of flavor and as collider targets [13, 14, 15]. Discrete symmetry approaches to structured Yukawa textures provide further motivation [16, 17].
Recent systematic scans of the Froggatt–Nielsen landscape [18, 19] have explored single-messenger models with integer charges, mapping their predictions onto the space of quark and lepton observables. Babu, Chandrasekar, and Tavartkiladze [20] have independently connected gauged flavor symmetries to the axion quality problem. These studies reinforce a key lesson: with a single flavon and integer FN exponents, the quark mass hierarchies can be accommodated, but the precision is limited by the coarseness of the integer lattice. The ninths-quantized $B$-lattice opens a qualitatively different region of model space—one in which multi-link chains of vectorlike fermions carrying discrete $\mathbb{Z}_{18}$ charges produce rational exponents, yielding substantially better fits to the observed mass and mixing patterns. Crucially, the same discrete gauge symmetry that enforces the lattice structure simultaneously protects the Peccei–Quinn symmetry against gravitational spoiling, as shown in the companion paper [6].
Unified Flavor (UF) unifies these ideas. A single flavon enforces the discrete $B$-lattice quantization of FN exponents, while TeV-scale nearest-neighbor VLQ chains generate the effective Yukawa couplings as path sums of allowed hops. This framework simultaneously:
  1. 1.
    generates hierarchical quark and lepton masses from a single parameter $B$;
  2. 2.
    produces CKM mixing structures consistent with the $B$-lattice flavor program [6, 7];
  3. 3.
    suppresses dangerous flavor-violating operators via chain locality;
  4. 4.
    requires TeV-scale VLQs accessible at the HL-LHC;
  5. 5.
    protects Peccei–Quinn axion quality through the same $\mathbb{Z}_{18}$ discrete gauge symmetry that enforces the lattice exponents, tying the flavor hierarchy to the strong CP problem.
Relation to earlier work. Papers I–IV of the $B$-lattice program established the lattice structure phenomenologically: Paper I [6] showed that a single parameter $B$ organizes quark and lepton masses; Paper II [5] introduced the two-over-two lattice exponent structure from three messenger chains; Paper III [7] derived analytic CKM phase relations from adjacent exponents; and Paper IV [8] extended the framework to lepton mixing and the PMNS matrix. The present work provides the dynamical ultraviolet (UV) realization: the lattice exponent algebra emerges from chain locality, rendering the hierarchy structural rather than merely parametric.
Table 1 summarizes the scope of each paper in the program.
Table 1: Roadmap of the $B$-lattice flavor program.
Paper Focus Ref.
I Single-$B$ mass organization [6]
II Two-over-two lattice exponents [5]
III CKM mixing & FX phase [7]
[…]
$Z_{18}$ axion quality [21]
V Chain UV completion (this work)
II Discrete Lattice Algebra
II.1 Symmetry structure
We impose the flavor symmetry
$G=Z_{N}\times Z_{2}^{\rm(NN)},$ (1)
where $Z_{N}$ (with $N=9$ or $18$) enforces the ninths-quantized suppression of FN operators, and $Z_{2}^{\rm(NN)}$ forbids nonlocal (non-nearest-neighbor) chain couplings. The discrete gauge anomaly cancellation conditions of Ibáñez and Ross [4] are satisfied generation by generation for $Z_{9}$ or $Z_{18}$.
A single flavon $\Phi$ carries charge
$q_{N}(\Phi)=1,$ (2)
and develops a VEV defining the expansion parameter
$\epsilon\equiv\frac{\langle\Phi\rangle}{\Lambda}=\frac{1}{B},\qquad B=\frac{75}{14}\simeq 5.357.$ (3)
II.2 Exponent lattice and algebraic closure
Allowed Yukawa exponents form the additive lattice
$\mathcal{L}=\left\{\frac{n}{9}\;\Big|\;n\in\mathbb{Z}\right\},$ (4)
so that if two operators carry exponents $a/9$ and $b/9$, their product carries $(a+b)/9$. Thus $\mathcal{L}$ is closed under composition. This algebraic closure ensures that the lattice structure is preserved under all allowed operator products—a property that neither continuous-$U(1)$ FN models nor generic clockwork constructions guarantee.
In the $B$-lattice program, the charged-lepton and quark sectors use a refined lattice with denominator $18$ (reflecting the bilinear structure of the Yukawa operator), while the neutrino Weinberg operator uses denominator $9$ [5]. The chain mechanism generates both denominators naturally: a single chain hop contributes $n/9$, while endpoint dressings involving both left- and right-handed fields contribute half-integer ninths, i.e. elements of $\mathcal{L}_{18}=\{n/18\}$.
III Field Content and Charge Assignments
III.1 Down-type vectorlike chain
A down-type VLQ chain is introduced with four sites:
$D_{a}+\bar{D}_{a},\qquad a=1,\dots,4,$ (5)
carrying $Z_{9}$ charges
$q_{9}(D_{a})=(0,\,8,\,6,\,2).$ (6)
The $Z_{2}^{\rm(NN)}$ symmetry assigns opposite parities to adjacent sites, forbidding non-nearest-neighbor mass terms. The chain structure is shown schematically in Fig. 1.
The allowed nearest-neighbor hops are
$\bar{D}_{1}\,D_{2}\;\Phi$ $:\quad\Delta q=0-8\equiv 1\!\!\pmod{9},\quad\text{suppression }\epsilon^{1/9},$ (7)
$\bar{D}_{2}\,D_{3}\;\Phi^{2}$ $:\quad\Delta q=8-6=2\!\!\pmod{9},\quad\text{suppression }\epsilon^{2/9},$ (8)
$\bar{D}_{3}\,D_{4}\;\Phi^{4}$ $:\quad\Delta q=6-2=4\!\!\pmod{9},\quad\text{suppression }\epsilon^{4/9},$ (9)
giving a discrete hop set $\{1,2,4\}/9$. Note that the number of flavon insertions at each hop equals the $Z_{9}$ charge difference, ensuring gauge invariance. The total chain suppression is $\epsilon^{(1+2+4)/9}=\epsilon^{7/9}$.
$D_{1}$$D_{2}$$D_{3}$$D_{4}$$q_{9}\!=\!0$$q_{9}\!=\!8$$q_{9}\!=\!6$$q_{9}\!=\!2$$\epsilon^{1/9}$$\epsilon^{2/9}$$\epsilon^{4/9}$$d_{R,j}$$q_{L,i}$$\epsilon^{B_{j}/9}$$\epsilon^{A_{i}/9}$$\Sigma=1+2+4=7$  (total hop exponent in ninths)
Figure 1: Schematic of the four-site down-type VLQ chain. Solid arrows denote nearest-neighbor hops with the indicated $\epsilon^{h_{a}/9}$ suppression; dashed arrows show the endpoint couplings to SM fields. The $Z_{9}$ charges are shown below each site. The effective Yukawa exponent for entry $(i,j)$ is $(\Sigma+A_{i}+B_{j})/9$, an exact result of the chain-inversion theorem (Sec. IV), not a perturbative approximation.
III.2 Endpoint couplings
The Standard Model (SM) quarks couple to the chain endpoints via
$\mathcal{L}_{\rm end}=\lambda_{i}\,\bar{q}_{L,i}\,\widetilde{H}\,D_{4}\,\left(\frac{\Phi}{\Lambda}\right)^{\!A_{i}}+\eta_{j}\,\bar{D}_{1}\,d_{R,j}\left(\frac{\Phi}{\Lambda}\right)^{\!B_{j}}+\text{h.c.},$ (10)
where $\widetilde{H}=i\sigma_{2}H^{*}$, and $A_{i}$, $B_{j}$ are integer or rational exponents fixed by the $Z_{N}$ charges of the SM fields $q_{L,i}$ and $d_{R,j}$. The coefficients $\lambda_{i}$ and $\eta_{j}$ are $\mathcal{O}(1)$ complex numbers.
III.3 Up-type chain
An analogous up-sector chain $U_{a}+\bar{U}_{a}$ ($a=1,\dots,N_{u}$) generates
$(Y_{u})_{ij}\sim\epsilon^{(\Sigma_{u}+A_{i}^{u}+B_{j}^{u})/9},$ (11)
where $\Sigma_{u}$ is the total hop suppression of the up chain. The two chains share the same left-handed doublet endpoints, as shown in Fig. 2. In a minimal realization, the CKM matrix receives contributions from both sectors:
$V_{\rm CKM}=U_{uL}^{\dagger}\,U_{dL},$ (12)
with $U_{uL}$ and $U_{dL}$ the left-handed rotations from diagonalizing $Y_{u}Y_{u}^{\dagger}$ and $Y_{d}Y_{d}^{\dagger}$ respectively. In the “down-dominant” limit where the up-sector texture is nearly diagonal, the CKM structure is controlled primarily by $U_{dL}$, and the analysis of Sec. VI applies directly. The full two-chain analysis, which includes CKM interference between different sectors, is presented in Appendix C.
[…]
Figure 2: Combined down-type (blue, upper) and up-type (red, lower) VLQ chains. Both chains share the same internal hop structure $\{1,2,4\}/9$ and the same left-handed endpoint dressings $A_{i}=(27,18,0)$, enforced by SU(2)L gauge invariance acting on the shared quark doublet $q_{L,i}$. The right-handed dressings differ: $B_{j}^{d}=(10,3,0)$ for the down sector and $B_{j}^{u}=(37,12,0)$ for the up sector. The $\epsilon^{h/9}$ suppressions and the resulting Yukawa exponents are exact consequences of the upper-triangular chain mass matrix (Sec. IV.5); they do not rely on $\kappa_{a}\epsilon^{h_{a}/9}\ll 1$. The CKM matrix arises from the mismatch between the left-handed rotations $U_{dL}$ and $U_{uL}$ that diagonalize the two Yukawa textures.
III.4 Why four chain sites
The number of chain sites $N=4$ is not a free choice but is determined by the requirement that the total chain hop exponent $\Sigma/9$ be large enough to generate the full hierarchy between the heaviest and lightest quarks of a given charge. For the down sector, the ratio $m_{d}/m_{b}\sim\epsilon^{37/9}$ requires a total suppression of $\epsilon^{37/9}\simeq 10^{-3}$. The large left-handed charge $Q(Q_{1})=3$ and right-handed charge $Q(d^{c}_{1})=10/9$ together provide the full hierarchy; the chain hops distribute these charges across nearest-neighbor links. A three-site chain with hop set $\{1,2\}/9$ would give $\Sigma=3$, yielding $\epsilon^{3/9}\simeq 0.57$—insufficient hierarchy. A five-site chain would provide more suppression than needed, at the cost of additional VLQ species. Four sites is thus the minimal choice consistent with the observed hierarchy and the discrete hop set $\{1,2,4\}/9$.
IV General Chain Inversion Theorem
The effective Yukawa coupling between SM quarks $q_{L,i}$ and $d_{R,j}$ is generated by integrating out the heavy VLQ chain that connects them. The SM fields do not couple directly to each other; $q_{L,i}$ couples to the last chain site $D_{N}$, and $d_{R,j}$ couples to the first site $D_{1}$ (Sec. III). The effective coupling is therefore proportional to the corner element $(\mathcal{M}^{-1})_{N1}$ of the inverse chain mass matrix—the amplitude for a signal injected at site $1$ to propagate through the chain and emerge at site $N$.
The $Z_{2}^{(\rm NN)}$ nearest-neighbor symmetry of Eq. (1) plays a crucial structural role here. By assigning opposite $Z_{2}$ parities to adjacent chain sites, it forbids all mass terms between non-adjacent sites: $\bar{D}_{a}D_{b}$ with |ab|>1|a-b|>1 is forbidden. In the $(D_{1},\dots,D_{N})$ basis, the chain mass matrix therefore has nonzero entries only on the diagonal (vectorlike masses $M_{a}$) and on the first superdiagonal (nearest-neighbor hops $m_{a,a+1}$), making it upper-triangular. This is not a simplifying approximation—it is an exact consequence of the symmetry. The payoff is that the inverse of an upper-triangular matrix is also upper-triangular and can be computed in closed form, so the corner element and hence the effective Yukawa exponent are exact, with no perturbative expansion required.
IV.1 $N$-site chain
Consider an $N$-site chain with upper-triangular mass matrix in the basis $(D_{1},\dots,D_{N})$:
$\mathcal{M}=\begin{pmatrix}M_{1}&m_{12}&0&\cdots&0\\ 0&M_{2}&m_{23}&\cdots&0\\ \vdots&&\ddots&\ddots&\vdots\\ 0&\cdots&0&M_{N-1}&m_{N-1,N}\\ 0&\cdots&0&0&M_{N}\end{pmatrix}$, (13)
where $M_{a}$ are the diagonal (vectorlike) masses and $m_{a,a+1}=\kappa_{a}M_{a}\,\epsilon^{h_{a}/9}$ are the nearest-neighbor hop masses, with $h_{a}$ the discrete hop exponents and $\kappa_{a}=\mathcal{O}(1)$.
IV.2 Block-inversion formula
Because $\mathcal{M}$ is upper-triangular, its inverse is also upper-triangular and can be computed in closed form by back-substitution [22] (see Appendix A for the explicit derivation). The corner element is exact:
$(\mathcal{M}^{-1})_{N1}=(-1)^{N-1}\frac{\prod_{k=1}^{N-1}m_{k,k+1}}{\prod_{a=1}^{N}M_{a}},$ (14)
with no series truncation and no requirement that $m_{a,a+1}\ll M_{a}$.
For degenerate masses $M_{a}=M$, this simplifies to
$\boxed{(\mathcal{M}^{-1})_{N1}=\frac{(-1)^{N-1}}{M}\prod_{k=1}^{N-1}\frac{m_{k,k+1}}{M}=\frac{(-1)^{N-1}\prod_{k}\kappa_{k}}{M}\,\epsilon^{\,\Sigma/9}}$ (15)
where $\Sigma\equiv\sum_{k=1}^{N-1}h_{k}$ is the total hop exponent and the overall sign is absorbed into the Yukawa coupling. Crucially, this formula holds for any values of $\kappa_{a}$ and $\epsilon^{h_{a}/9}$—it is an algebraic identity, not the leading term of a perturbative expansion.
IV.3 Explicit $4$-site example
For the down chain of Sec. III, the mass matrix is
$\mathcal{M}_{D}=\begin{pmatrix}M&\kappa_{1}M\epsilon^{1/9}&0&0\\ 0&M&\kappa_{2}M\epsilon^{2/9}&0\\ 0&0&M&\kappa_{3}M\epsilon^{4/9}\\ 0&0&0&M\end{pmatrix}$. (16)
The block-inversion formula (15) gives
$(\mathcal{M}_{D}^{-1})_{41}=-\frac{\kappa_{1}\kappa_{2}\kappa_{3}}{M}\,\epsilon^{(1+2+4)/9}=-\frac{\kappa_{1}\kappa_{2}\kappa_{3}}{M}\,\epsilon^{7/9},$ (17)
where the sign is absorbed into the Yukawa coupling. This result is exact: numerical evaluation for $\kappa_{a}=1$ gives $|(\mathcal{M}_{D}^{-1})_{41}|=\epsilon^{7/9}/M$ to machine precision, and the agreement persists even for $\kappa_{a}\gg 1$.
IV.4 Path-Sum Theorem
Combining the chain corner element (17) with the endpoint couplings (10), the effective Yukawa matrix is
$(Y_{d})_{ij}=\lambda_{i}\eta_{j}\,\frac{\kappa_{1}\kappa_{2}\kappa_{3}}{M}\,\epsilon^{(\Sigma+A_{i}+B_{j})/9},$ (18)
with $\Sigma=7$.
Path-Sum Theorem. In a nearest-neighbor lattice chain, the effective Yukawa exponent for each entry $(i,j)$ equals the sum of the discrete hop exponents $\Sigma=\sum h_{k}$ plus the endpoint dressings $A_{i}$ and $B_{j}$: $p_{ij}=\frac{\Sigma+A_{i}+B_{j}}{9}.$ (19) This realizes the $B$-lattice exponent algebra dynamically.
IV.5 Convergence of the chain expansion
A potential concern is that the individual hop parameters $\kappa_{a}\,\epsilon^{h_{a}/9}$ are not uniformly small—the smallest hop has $h_{1}=1$, giving $\epsilon^{1/9}\simeq 0.83$, so that for $\kappa_{1}\sim 1$ the ratio $m_{12}/M\sim 0.83$ is not a conventional small parameter. We now show that this does not compromise the framework, because the quantity $\kappa_{a}\epsilon^{h_{a}/9}$ is never used as a perturbative expansion parameter.
The chain inversion is exact.
The corner element formula (14)–(15) is an algebraic identity for upper-triangular matrices, derived by back-substitution with no series truncation (see Appendix A). As verified numerically, the formula reproduces the exact matrix inverse to machine precision for all values of $\kappa_{a}$, including $\kappa_{a}\,\epsilon^{h_{a}/9}\gg 1$. The $Z_{2}^{(\rm NN)}$ locality symmetry enforces the upper-triangular structure, and the Path-Sum Theorem therefore holds without any small-parameter assumption.
The Yukawa diagonalization uses different expansion parameters.
The perturbative diagonalization of the effective $3\times 3$ Yukawa matrix (Sec. VI and Appendix C) expands in the left-handed mixing angles $\theta_{ij}^{dL}\sim\epsilon^{|Q(Q_{i})-Q(Q_{j})|}$, whose magnitudes are
$\theta_{12}^{dL}\sim\epsilon\simeq 0.19,\quad\theta_{23}^{dL}\sim\epsilon^{2}\simeq 0.035,\quad\theta_{13}^{dL}\sim\epsilon^{3}\simeq 0.007.$ (20)
These are the physical expansion parameters of the diagonalization, and all are comfortably small. The quantity $\epsilon^{1/9}\simeq 0.83$ does not appear as an expansion parameter at this stage.
What $\kappa_{a}\epsilon^{h_{a}/9}\sim\mathcal{O}(1)$ does affect.
The chain mass eigenstates receive $\mathcal{O}(\kappa^{2}\epsilon^{2h_{a}/9})$ shifts relative to the site-basis states (see Sec. IX), and these splittings can be large: $\sim 69\%$ for the first hop ($h_{1}=1$) and $\sim 47\%$ for the second ($h_{2}=2$). These corrections redistribute mass among the VLQ eigenstates but do not alter the exponent structure of the effective Yukawa. The VLQ spectrum and SM–VLQ mixing angles (Sec. VIII) should therefore be understood as order-of-magnitude estimates, with $\mathcal{O}(1)$ corrections from the non-perturbative regime of the chain splittings.
Robustness under $\mathcal{O}(1)$ variation.
Exact numerical diagonalization of the $3\times 3$ down-type Yukawa for $10^{3}$ random $\mathcal{O}(1)$ complex coefficients $c^{d}_{ij}$ with $|c^{d}_{ij}|\in[0.5,\,2.0]$ confirms that the eigenvalue ratios $y_{d}/\epsilon^{37/9}$, $y_{s}/\epsilon^{7/3}$, and $y_{b}/1$ remain $\mathcal{O}(1)$ in every case, with means of $2.2$, $1.6$, and $1.5$ respectively. The $\epsilon$-scaling is therefore robust: higher-order corrections modify only the $\mathcal{O}(1)$ prefactors.
V Connection to the $B$-Lattice Quark Textures
The $B$-lattice quark program [6, 5, 7] established the Yukawa exponent matrices
$p^{u}_{ij}$ $={\begin{pmatrix}\tfrac{64}{9}&\tfrac{13}{3}&3\\[2.0pt] \tfrac{55}{9}&\tfrac{10}{3}&2\\[2.0pt] \tfrac{37}{9}&\tfrac{4}{3}&0\end{pmatrix}}$, (21)
$p^{d}_{ij}$ $={\begin{pmatrix}\tfrac{37}{9}&\tfrac{10}{3}&3\\[2.0pt] \tfrac{28}{9}&\tfrac{7}{3}&2\\[2.0pt] \tfrac{10}{9}&\tfrac{1}{3}&0\end{pmatrix}}$.
These factorize as $p^{f}_{ij}=Q(Q_{i})+Q(f^{c}_{j})$ with shared left-handed doublet charges $Q(Q_{i})=(3,2,0)$ and right-handed charges $Q(u^{c}_{j})=(\tfrac{37}{9},\tfrac{4}{3},0)$, $Q(d^{c}_{j})=(\tfrac{10}{9},\tfrac{1}{3},0)$.
These charges are not free parameters chosen to fit the data; they are determined bottom-up from the observed mass spectrum in a three-step sequence that leaves no free choices.
First, the value $B=75/14$ is itself fixed by the requirement that the six quark mass ratios $m_{f}/m_{f_{3}}=\epsilon^{p_{f}}$ simultaneously yield $\mathcal{O}(1)$ coefficients $c_{f}\simeq 1.00$ [6]. No other value of $B$ reproduces all six ratios to this precision (Sec. V.1).
Second, once the diagonal exponents $n_{f}$ are read off from the measured quark masses at $M_{Z}$, the factorization $p^{f}_{ij}=Q(Q_{i})+Q(f^{c}_{j})$ is an overdetermined system: $9$ entries per sector are constrained by $3+3$ charges, so $12$ independent exponents must accommodate $18$ texture entries across both sectors. With the convention $Q(Q_{3})=Q(f^{c}_{3})=0$ (third generation carries zero charge), the decomposition is unique: $Q(Q_{i})=p^{d}_{i3}$ and $Q(f^{c}_{j})=p^{f}_{3j}$, read directly from the third row and column of the exponent matrix. That the resulting charges are rational multiples of $1/9$ is a consequence of the observed mass ratios, not a design choice.
Third, the factorization succeeds—with all $\mathcal{O}(1)$ coefficients essentially unity (Sec. V.1)—which is a nontrivial check, not a parametric fit. The subsequent verification that these phenomenologically determined charges satisfy the discrete anomaly conditions (Sec. XIII) is therefore a prediction of the framework, not an input.
The Yukawa matrices are
$(Y_{f})_{ij}=c^{f}_{ij}\,\epsilon^{\,p^{f}_{ij}},\qquad f=u,\,d,$ (22)
where the $c^{f}_{ij}$ are $\mathcal{O}(1)$ complex coefficients encoding the product of vertex couplings along the chain.
In the multi-messenger realization of Ref. [5], each Yukawa entry receives contributions from several chain configurations (messenger species), so that the effective coefficient takes the form
$C^{f}_{ij}=1+e^{i\phi_{f}}\,\epsilon^{\Delta^{f}_{ij}}+e^{i\psi_{f}}\,\epsilon^{\Delta^{\prime f}_{ij}},$ (23)
where $\Delta^{f}_{ij},\Delta^{\prime f}_{ij}\geq 0$ are additional suppression exponents from the subleading chains, and $\phi_{f}$, $\psi_{f}$ are relative phases. The terminology “multi-messenger” is borrowed, with intent, from astrophysics: just as multi-messenger astronomy combines independent observation channels—gravitational waves, neutrinos, photons—to reconstruct a single astrophysical event, multi-messenger flavor physics combines independent heavy-fermion chains to construct each Yukawa coupling, with the interference pattern encoding the CP-violating phase. The leading term is unity, and the corrections are suppressed by positive powers of $\epsilon$, so $|C^{f}_{ij}|$ is generically $\mathcal{O}(1)$ without fine tuning. The phases $\phi_{f}$ and $\psi_{f}$ are the physical origin of CP violation: they generate the CKM phase through the adjacent-exponent interference mechanism discussed in Sec. VII. The $\mathcal{O}(1)$ coefficients $c^{f}_{ij}$ in Eq. (22) are identified with these effective coefficients, $c^{f}_{ij}=C^{f}_{ij}\times(\text{overall normalization})$.
For $\epsilon\ll 1$, the hierarchy is controlled by the exponent matrix $p^{f}$; the coefficients provide $\mathcal{O}(1)$ modulation but do not alter the power-counting.
Each Yukawa matrix is diagonalized by a bidiagonal (singular-value) decomposition:
$U_{fL}^{\dagger}\,Y_{f}\,U_{fR}=Y_{f}^{\rm diag}\equiv\mathrm{diag}(y_{f_{1}},\,y_{f_{2}},\,y_{f_{3}}),$ (24)
where $U_{fL}$ and $U_{fR}$ are unitary matrices acting on the left-handed and right-handed fields respectively, and the $y_{f_{i}}$ are the real, non-negative Yukawa eigenvalues related to the quark masses by $m_{f_{i}}=y_{f_{i}}\,v/\sqrt{2}$. The CKM matrix is then
$V_{\rm CKM}=U_{uL}^{\dagger}\,U_{dL}.$ (25)
The quark mass hierarchies are
$m_{u}:m_{c}:m_{t}\sim\epsilon^{64/9}:\epsilon^{10/3}:1,$ (26)
$m_{d}:m_{s}:m_{b}\sim\epsilon^{37/9}:\epsilon^{7/3}:1,$ (27)
and the CKM mixing angles scale as
$|V_{us}|$ $\sim\epsilon^{8/9}\simeq 0.22,$ (28)
$|V_{cb}|$ $\sim\epsilon^{17/9}\simeq 0.042,$ (29)
$|V_{ub}|$ $\sim\epsilon^{10/3}\simeq 0.0035,$ (30)
in excellent agreement with the Particle Data Group (PDG) values [23].
In the chain framework, these exponents are reproduced by the path-sum formula
$p^{f}_{ij}=\frac{A_{i}+B_{j}^{f}}{9},$ (31)
where $A_{i}$ are the shared left-handed endpoint dressings (in ninths) and $B_{j}^{f}$ are the right-handed dressings. SU(2)L gauge invariance requires the same $A_{i}$ for both sectors:
$A_{i}=(27,\,18,\,0),$ (32)
with
$B_{j}^{d}=(10,\,3,\,0),\qquad B_{j}^{u}=(37,\,12,\,0).$ (33)
The chain hop contributions are absorbed into the endpoint dressings, so the factorized charge structure $p^{f}_{ij}=Q(Q_{i})+Q(f^{c}_{j})$ is manifest. This demonstrates that the phenomenologically successful $B$-lattice textures emerge naturally from the discrete charge assignments of Sec. III.
V.1 Numerical benchmark
With $\epsilon=14/75\simeq 0.1867$ and $\overline{\rm MS}$ quark masses at $M_{Z}$ [6, 23] ($m_{u}=1.11$ MeV, $m_{d}=2.82$ MeV, $m_{s}=55.7$ MeV, $m_{c}=0.629$ GeV, $m_{b}=2.794$ GeV, $m_{t}=169$ GeV), the $B$-lattice gives:
$\frac{m_{d}}{m_{b}}\simeq\epsilon^{37/9}\cdot c_{d}=1.01\times 10^{-3}\cdot c_{d},$ (34)
$\frac{m_{s}}{m_{b}}\simeq\epsilon^{7/3}\cdot c_{s}=1.99\times 10^{-2}\cdot c_{s},$ (35)
$\frac{m_{u}}{m_{t}}\simeq\epsilon^{64/9}\cdot c_{u}=6.6\times 10^{-6}\cdot c_{u},$ (36)
$\frac{m_{c}}{m_{t}}\simeq\epsilon^{10/3}\cdot c_{c}=3.72\times 10^{-3}\cdot c_{c},$ (37)
with $c_{d}\simeq c_{s}\simeq c_{u}\simeq c_{c}\simeq 1.00$. All four $\mathcal{O}(1)$ coefficients are essentially unity—the $B$-lattice exponents of Eq. (21) reproduce the physical quark mass ratios $m_{d}/m_{b}\simeq 1.01\times 10^{-3}$, $m_{s}/m_{b}\simeq 1.99\times 10^{-2}$, $m_{u}/m_{t}\simeq 6.6\times 10^{-6}$, and $m_{c}/m_{t}\simeq 3.7\times 10^{-3}$ with no fine tuning.
VI Perturbative Diagonalization of the Down-Type Texture
We illustrate the sequential diagonalization procedure using the physical down-type texture of Eq. (21). Setting all $\mathcal{O}(1)$ coefficients to unity for transparency, the Yukawa matrix is
$Y_{d}\sim\begin{pmatrix}\epsilon^{37/9}&\epsilon^{10/3}&\epsilon^{3}\\[3.0pt] \epsilon^{28/9}&\epsilon^{7/3}&\epsilon^{2}\\[3.0pt] \epsilon^{10/9}&\epsilon^{1/3}&1\end{pmatrix}$. (38)
Because $p^{d}_{ij}=Q(Q_{i})+Q(d^{c}_{j})$ factorizes, the matrix is not symmetric: the left-handed charges $Q(Q_{i})=(3,2,0)$ differ from the right-handed charges $Q(d^{c}_{j})=(\tfrac{10}{9},\tfrac{1}{3},0)$. The diagonalization therefore requires separate left- and right-handed rotations, $U_{dL}^{\dagger}\,Y_{d}\,U_{dR}=Y_{d}^{\rm diag}$.
VI.1 Third generation
The $(3,3)$ element dominates the matrix ($p^{d}_{33}=0$, i.e. unsuppressed):
$y_{b}\simeq 1.$ (39)
The physical $b$-quark mass is set by the overall Yukawa normalization $m_{b}=y_{b}\,v/\sqrt{2}$. In the companion analysis of Ref. [6], this is anchored to the charged-lepton sector through $m_{b}^{\rm ref}=\varphi\,m_{\tau}^{\overline{\rm MS}}(M_{Z})$, where $\varphi=(1+\sqrt{5})/2$ is the golden ratio—a Georgi–Jarlskog-style relation that fixes the reference scale from which all other quark masses follow via $m_{f}=m_{b}^{\rm ref}\cdot B^{\,n_{f}}$.
VI.2 2233 rotation
The $(2,3)$ element is removed by a left-handed rotation with angle
$\theta_{23}^{dL}\approx\frac{|(Y_{d})_{23}|}{|(Y_{d})_{33}|}=\frac{\epsilon^{2}}{1}=\epsilon^{2}\simeq 0.035.$ (40)
This is parametrically $\epsilon^{Q(Q_{2})-Q(Q_{3})}=\epsilon^{2}$. The right-handed rotation has a different angle, $\theta_{23}^{dR}\approx\epsilon^{1/3}\simeq 0.57$, reflecting the asymmetry between left and right charges.
VI.3 Second generation
After decoupling the third generation, the effective $(2,2)$ element is
$y_{s}\simeq\epsilon^{7/3}-\frac{\epsilon^{2}\cdot\epsilon^{1/3}}{1}=\epsilon^{7/3}(1-\epsilon^{0})+\cdots,$ (41)
which at leading order gives
$y_{s}\sim\epsilon^{7/3}\simeq 0.020,$ (42)
since the seesaw correction from the 2233 block involves $\epsilon^{2+1/3}=\epsilon^{7/3}$, the same power as the diagonal element itself. The physical coefficient thus depends on the $\mathcal{O}(1)$ factors; generically $y_{s}=|c_{22}^{d}-c_{23}^{d}c_{32}^{d}/c_{33}^{d}|\,\epsilon^{7/3}$.
VI.4 1122 rotation
The residual $(1,2)$ element after the 2233 rotation scales as $\epsilon^{10/3}$. The 1122 left-handed mixing angle is therefore
$\theta_{12}^{dL}\approx\frac{\epsilon^{10/3}}{\epsilon^{7/3}}=\epsilon^{10/3-7/3}=\epsilon\simeq 0.187.$ (43)
This is $\epsilon^{Q(Q_{1})-Q(Q_{2})}=\epsilon^{3-2}=\epsilon$.
VI.5 First generation and 1133 mixing
The lightest eigenvalue emerges as
$y_{d}\sim\epsilon^{37/9}\simeq 1.0\times 10^{-3}.$ (44)
The direct 1133 mixing angle is
$\theta_{13}^{dL}\approx\frac{\epsilon^{3}}{1}=\epsilon^{3}\simeq 0.007.$ (45)
This is $\epsilon^{Q(Q_{1})-Q(Q_{3})}=\epsilon^{3}$.
VI.6 Summary of eigenvalue and mixing hierarchy
Collecting the results for the down sector:
$y_{b}$ $\sim 1,$ $y_{s}$ $\sim\epsilon^{7/3},$ $y_{d}$ $\sim\epsilon^{37/9},$ (46)
$\theta_{23}^{dL}$ $\sim\epsilon^{2},$ $\theta_{12}^{dL}$ $\sim\epsilon,$ $\theta_{13}^{dL}$ $\sim\epsilon^{3}.$ (47)
Each left-handed mixing angle is controlled by the doublet charge difference $\theta_{ij}^{dL}\sim\epsilon^{|Q(Q_{i})-Q(Q_{j})|}$.
The right-handed rotation matrix $U_{dR}$ is determined by the right-handed charge differences $Q(d^{c}_{i})-Q(d^{c}_{j})$:
$\theta_{23}^{dR}$ $\sim\epsilon^{1/3}\simeq 0.57,$ $\theta_{12}^{dR}$ $\sim\epsilon^{7/9}\simeq 0.27,$
$\theta_{13}^{dR}$ $\sim\epsilon^{10/9}\simeq 0.15.$ (48)
Since the down-type right-handed charges $Q(d^{c}_{j})=(\tfrac{10}{9},\tfrac{1}{3},0)$ are more closely spaced than the left-handed charges $Q(Q_{i})=(3,2,0)$, the right-handed rotations are generically larger than the left-handed ones. The right-handed rotations do not enter the CKM matrix but are relevant for right-handed currents in extensions beyond the SM.
For the up sector, the same logic gives $\theta_{ij}^{uL}\sim\theta_{ij}^{dL}$ (shared doublet charges), while the right-handed angles are
$\theta_{23}^{uR}$ $\sim\epsilon^{4/3}\simeq 0.11,$ $\theta_{12}^{uR}$ $\sim\epsilon^{25/9}\simeq 0.009,$
$\theta_{13}^{uR}$ $\sim\epsilon^{37/9}\simeq 0.001.$ (49)
The up-type right-handed charges $Q(u^{c}_{j})=(\tfrac{37}{9},\tfrac{4}{3},0)$ are more widely spaced, yielding smaller (more hierarchical) right-handed rotations than in the down sector. Because the up sector shares the same doublet charges $Q(Q_{i})=(3,2,0)$, the up-type left-handed angles have identical $\epsilon$-scaling: $\theta_{ij}^{uL}\sim\theta_{ij}^{dL}$. The CKM elements therefore arise from the interference $V_{ij}\sim\theta_{ij}^{dL}-\theta_{ij}^{uL}\,e^{i\phi}$, as derived in Appendix C.
A key feature of the non-symmetric texture is that $\theta_{12}^{dL}\sim\epsilon\neq\theta_{23}^{dL}\sim\epsilon^{2}$, so the physical hierarchy $|V_{us}|\gg|V_{cb}|$ is built into the charge assignments. This resolves the degeneracy $\theta_{12}\sim\theta_{23}$ that would appear in a simplified symmetric texture with a common charge difference.
VI.7 Jarlskog invariant scaling
The Jarlskog invariant [24] scales as
$J\sim\theta_{12}\,\theta_{23}\,\theta_{13}\,\sin\delta.$ (50)
Using the full $B$-lattice scalings (28)–(30):
$J\sim\epsilon^{8/9}\cdot\epsilon^{17/9}\cdot\epsilon^{10/3}\cdot\sin\delta\sim\epsilon^{19/3}\,\sin\delta\simeq 2.5\times 10^{-5}\sin\delta,$ (51)
consistent with the PDG value $J=(3.00\pm 0.13)\times 10^{-5}$ [23] for $\sin\delta\sim\mathcal{O}(1)$.
VII CKM Phase from Adjacent Exponent Interference
Physical CP violation arises from interference between Yukawa entries whose exponents differ by $\mathcal{O}(1/9)$. To see this, introduce a single complex endpoint coefficient in the 2233 entry:
$Y_{23}=c_{1}\,\epsilon^{4/9}+c_{2}\,\epsilon^{5/9}\,e^{i\alpha},$ (52)
where $c_{1},c_{2}=\mathcal{O}(1)$ are real and $\alpha$ is a physical phase originating from the complex flavon VEV. The CKM phase is then
$\delta_{\rm CKM}\simeq\arg\!\left(1+r\,\epsilon^{1/9}\,e^{i\alpha}\right),$ (53)
with $r=c_{2}/c_{1}$. Since $\epsilon^{1/9}\simeq 0.83$, the interference is $\mathcal{O}(1)$ and generates a large CP phase without fine tuning. (Note that $\epsilon^{1/9}$ being $\mathcal{O}(1)$ is a virtue for CP violation; as shown in Sec. IV.5, it does not compromise the chain inversion or the Yukawa diagonalization.) This is the same mechanism identified in Paper III [7] and is a direct consequence of the densely spaced ninths lattice.
VII.1 Structural down-dominance
A natural question is whether the CP-violating phase requires a specific relationship between the up- and down-sector multi-messenger phases $\phi_{f}$, $\psi_{f}$. We now show that the CKM phase is structurally controlled by the down sector, so that the up-sector phases are largely irrelevant.
The Fritzsch–Xing decomposition (Appendix C) expresses the Cabibbo angle as
$|V_{us}|\simeq|s_{d}-s_{u}\,e^{i\phi_{\rm FX}}|,$ (54)
with $s_{d}\sim\epsilon^{8/9}\simeq 0.22$ and $s_{u}\sim\epsilon^{17/9}\simeq 0.042$. The ratio
$\frac{s_{u}}{s_{d}}=\epsilon^{17/9-8/9}=\epsilon\simeq 0.19$ (55)
quantifies the structural suppression: the up-sector contribution enters $|V_{us}|$ at the $\sim\!19\%$ level, regardless of the FX phase $\phi_{\rm FX}$. Varying $\phi_{\rm FX}$ over its full range changes $|V_{us}|$ by only $\pm s_{u}\simeq 0.04$ around the down-sector baseline $s_{d}\simeq 0.22$. For $|V_{cb}|$ the suppression is even stronger: the up-sector correction enters at order $\epsilon^{17/18}\simeq 0.20$ relative to the down-sector contribution.
The physical origin of this hierarchy is the asymmetry in right-handed charge spacings:
$Q(d^{c}_{j})=\bigl(\tfrac{10}{9},\;\tfrac{1}{3},\;0\bigr)\quad\text{vs.}\quad Q(u^{c}_{j})=\bigl(\tfrac{37}{9},\;\tfrac{4}{3},\;0\bigr).$ (56)
The wider spacing of the up-sector charges produces a more hierarchical up-type Yukawa, so that the up-sector eigenvalue ratios $m_{u}/m_{c}$ and $m_{c}/m_{t}$ are much smaller than their down-sector counterparts $m_{d}/m_{s}$ and $m_{s}/m_{b}$. The CKM mixing angles—being differences of left-handed rotations from the two sectors—are therefore dominated by whichever sector has the larger eigenvalue ratios, which is structurally the down sector.
The benchmark phases of Appendix D (vanishing $\phi_{u}$, $\psi_{u}$; near-$\pi$ values of $\phi_{d}$, $\psi_{d}$) are therefore an illustrative simplification, not a fine-tuned choice. The vanishing up-sector phases can be replaced by arbitrary $\mathcal{O}(1)$ values with only a $\sim\!19\%$ change in $|V_{us}|$ and a correspondingly small shift in the Jarlskog invariant.
VII.2 Genericity of $\mathcal{O}(1)$ CP violation
To quantify the extent to which large CP violation is a generic prediction of the lattice structure (rather than a consequence of specific phase choices), we perform a numerical scan over random $\mathcal{O}(1)$ complex coefficients $c^{f}_{ij}$ with $|c^{f}_{ij}|\in[0.5,\,2.0]$ and uniformly distributed phases, independently for both sectors. For each of $5\times 10^{3}$ trials, the full Yukawa matrices (22) are constructed and the CKM matrix is extracted by exact singular-value decomposition.
The results (Table 2) demonstrate that:
  1. 1.
    The Jarlskog invariant falls within a factor of 3 of the PDG value $J=(3.00\pm 0.13)\times 10^{-5}$ [23] in $55\%$ of all trials, and within the range $10^{-6},\,10^{-3}]$ in $98\%$ of trials.
  2. 2.
    The effective $|\!\sin\delta|$ exceeds $0.3$ in $81\%$ of trials—large CP violation is the generic outcome.
  3. 3.
    The CKM mixing magnitudes $|V_{us}|$, $|V_{cb}|$, and $|V_{ub}|$ cluster around their PDG values with $\mathcal{O}(1)$ scatter.
Table 2: CKM observables from $5\times 10^{3}$ random $\mathcal{O}(1)$ complex coefficient trials. “Median” and “IQR” denote the median and interquartile range.
Observable Median IQR PDG
$J/10^{-5}$ 4.1 [1.2, 10][1.2,\;10] $3.0\pm 0.1$
$|V_{us}|$ 0.26 [0.16, 0.39][0.16,\;0.39] 0.225
$|V_{cb}|$ 0.042 [0.025, 0.060][0.025,\;0.060] 0.042
$|V_{ub}|$ 0.010 [0.004, 0.017][0.004,\;0.017] 0.004
Phase counting. The standard phase-counting theorem for three generations gives (n1)(n2)/2=1(n-1)(n-2)/2=1 physical CP-violating phase [24]. The multi-messenger phases $\phi_{f}$, $\psi_{f}$ are not additional parameters: they are specific combinations of the UV coupling phases that survive after all field rephasing. The chain framework therefore has the same number of irreducible phase inputs as the SM—one. Its advantage is that the lattice exponent algebra determines the scaling $J\sim\epsilon^{19/3}\sin\delta$ structurally, while the dense ninths spacing ensures that the remaining free parameter $\sin\delta$ is generically $\mathcal{O}(1)$.
VIII Flavor Constraints: $\Delta F=2$ and FCNCs
VIII.1 SM–VLQ mixing
Integrating out the heavy chain states induces mixing between SM quarks and VLQs. The left-handed mixing angles scale as
$\theta_{L}^{i}\sim\lambda_{i}\,\frac{v}{M}\,\epsilon^{n_{i}/9},$ (57)
where $v\simeq 174$ GeV is the electroweak VEV, $M$ is the common VLQ mass, $n_{i}$ encodes the total suppression for generation $i$, and $\lambda_{i}$ is the $\mathcal{O}(1)$ endpoint coupling from Eq. (10).
For a benchmark mass $M=2$ TeV, $\lambda_{i}\sim 1$, and the $B$-lattice charges, the angles are
$\theta_{L}^{b}$ $\sim\lambda_{3}\,\frac{v}{M}\,\epsilon^{0}\sim 10^{-1}\,\lambda_{3},$ (58)
$\theta_{L}^{s}$ $\sim\lambda_{2}\,\frac{v}{M}\,\epsilon^{7/3}\sim 2\times 10^{-3}\,\lambda_{2},$ (59)
$\theta_{L}^{d}$ $\sim\lambda_{1}\,\frac{v}{M}\,\epsilon^{37/9}\sim 10^{-4}\,\lambda_{1}.$ (60)
The precise value of $\lambda_{3}$ is constrained by the $Zb\bar{b}$ vertex measurement (see Sec. VIII.4); all other angles carry additional $\epsilon$-suppression from the lattice hierarchy.
VIII.2 $\Delta F=2$ operators
The dominant new-physics contribution to meson mixing arises from tree-level VLQ exchange, generating four-fermion operators suppressed by
$C_{\Delta F=2}\sim\frac{(\theta_{L}^{i})^{2}(\theta_{L}^{j})^{2}}{M^{2}}.$ (61)
For $K¯\bar{K}$ mixing ($i=d$, $j=s$):
$C_{K\bar{K}}\sim\frac{(10^{-4})^{2}(2\times 10^{-3})^{2}}{(2~\text{TeV})^{2}}\sim 10^{-18}~\text{GeV}^{-2},$ (62)
well below the experimental bound $C_{K}^{\rm exp}\lesssim 10^{-12}~\text{GeV}^{-2}$ [25].
For $B_{s}$$\bar{B}_{s}$ mixing ($i=s$, $j=b$):
$C_{B_{s}}\sim\frac{(2\times 10^{-3})^{2}(10^{-1})^{2}}{(2~\text{TeV})^{2}}\sim 10^{-14}~\text{GeV}^{-2},$ (63)
also consistent with current constraints.
The key point is that the $B$-lattice hierarchy in the mixing angles (58)–(60) provides a built-in Glashow–Iliopoulos–Maiani (GIM)-like suppression [26] of FCNCs—the same lattice that generates the mass hierarchy simultaneously protects against flavor violation.
VIII.3 One-loop dipole operators
The electromagnetic dipole Wilson coefficient $C_{7}$, which governs $b\to s\gamma$, receives three types of VLQ contributions:
(i) Modified $Wtb$ vertex. The SM–VLQ mixing shifts the $Wtb$ coupling by $\delta g_{Wtb}/g_{Wtb}\sim-(\theta_{L}^{b})^{2}/2$, modifying the dominant SM one-loop diagram. The fractional shift in $C_{7}$ is
$\frac{\delta C_{7}}{C_{7}^{\rm SM}}\sim-(\theta_{L}^{b})^{2}\sim-3\times 10^{-3},$ (64)
two orders of magnitude below the experimental sensitivity $|\delta C_{7}/C_{7}^{\rm SM}|\lesssim 0.10$ from the measured ${\rm BR}(B\to X_{s}\gamma)$ [23].
(ii) WW–VLQ loop. A new penguin diagram with a VLQ running in the loop alongside the $W$ boson contributes
$C_{7}^{\rm VLQ}\sim\frac{\alpha_{W}}{4\pi}\,\frac{\theta_{L}^{b}\,\theta_{L}^{s}}{V_{tb}\,V_{ts}}\,f_{7}\!\left(\frac{m_{t}^{2}}{M^{2}}\right),$ (65)
where $f_{7}(x)\to 7x/36$ for $x\ll 1$ and $\alpha_{W}=g^{2}/(4\pi)\simeq 0.034$. For $M=2$ TeV this is $C_{7}^{\rm VLQ}\sim 10^{-8}$, negligible.
(iii) $Z$-FCNC penguin. The tree-level $Z$-mediated $b\text{--}s$ FCNC coupling (see below) generates a one-loop dipole through a ZZtt penguin. This carries an additional loop suppression $\alpha_{\rm em}/(4\pi)\sim 6\times 10^{-4}$ beyond the tree-level $Z$-FCNC and is completely negligible: $C_{7}^{Z\text{-FCNC}}\sim 10^{-5}$.
The chromomagnetic dipole $C_{8}$ has the same parametric scaling as $C_{7}$ and is equally safe. In summary, all one-loop dipole contributions satisfy $|\delta C_{7}/C_{7}^{\rm SM}|\lesssim 10^{-2}$, well below current and projected experimental precision.
VIII.4 $Z$-mediated rare decays and $R_{b}$
Integrating out the SU(2)-singlet VLQ chain states at tree level generates a flavor-changing $Z$ coupling
$\delta g_{L}^{bs}=\frac{g}{2c_{W}}\;\frac{1}{2}\;\theta_{L}^{b}\,\theta_{L}^{s},$ (66)
where the factor $\tfrac{1}{2}=|I_{3}^{d_{L}}-I_{3}^{D}|$ reflects the isospin mismatch between the SM doublet and the SU(2)-singlet VLQ.
Semileptonic operators $C_{9}$, $C_{10}$.
Tree-level $Z$ exchange between the $b\text{--}s$ FCNC vertex and the lepton pair generates contributions to the Wilson coefficients $C_{9}$ (vector) and $C_{10}$ (axial):
$C_{9}^{\rm NP}$ $\sim\frac{\pi}{\alpha_{\rm em}}\,\frac{\delta g_{L}^{bs}\,(\!-\tfrac{1}{2}+2s_{W}^{2})}{\sqrt{2}\,G_{F}\,V_{tb}\,V_{ts}\,M_{Z}^{2}}\,$, (67)
$C_{10}^{\rm NP}$ $\sim-\frac{\pi}{\alpha_{\rm em}}\,\frac{\delta g_{L}^{bs}\times\tfrac{1}{2}}{\sqrt{2}\,G_{F}\,V_{tb}\,V_{ts}\,M_{Z}^{2}}\,.$ (68)
Because the $Z$-lepton vector coupling $(-\tfrac{1}{2}+2s_{W}^{2})\simeq-0.04$ is accidentally small, $C_{9}^{\rm NP}$ is suppressed relative to $C_{10}^{\rm NP}$ by a factor of $\sim 2|{-}\tfrac{1}{2}+2s_{W}^{2}|\simeq 0.08$. $C_{10}$ therefore provides the dominant constraint from rare $B$ decays.
$B_{s}\to\mu^{+}\mu^{-}$.
The branching ratio scales as $\mathrm{BR}\propto|C_{10}^{\rm SM}+C_{10}^{\rm NP}|^{2}$. The current measurement $\mathrm{BR}(B_{s}\to\mu^{+}\mu^{-})=(3.34\pm 0.27)\times 10^{-9}$ [23] is consistent with the SM prediction to $\sim\!15\%$ precision, requiring $|C_{10}^{\rm NP}|\lesssim 0.6$.
$R_{b}$ vertex correction.
The same SM–VLQ mixing that generates the FCNC also shifts the diagonal $Zb\bar{b}$ coupling:
$\delta g_{L}^{bb}=\frac{1}{2}(\theta_{L}^{b})^{2},$ (69)
producing a fractional change $\delta R_{b}/R_{b}\simeq-2\,|g_{L}^{b}|\,\delta g_{L}^{bb}/(g_{L}^{b\,2}+g_{R}^{b\,2})\simeq-2.3\,(\theta_{L}^{b})^{2}$. The LEP measurement $R_{b}=0.21629\pm 0.00066$ [23] constrains $|\delta R_{b}/R_{b}|\lesssim 0.3\%$, yielding the bound
$\theta_{L}^{b}\lesssim 0.04.$ (70)
This is the strongest individual constraint on $\theta_{L}^{b}$ in the chain framework, tighter than the oblique parameters.
Consistency.
The mixing angle was estimated in Eq. (58) as $\theta_{L}^{b}\sim\lambda_{3}\,v/M$, where $\lambda_{3}$ is the $\mathcal{O}(1)$ endpoint coupling of the third-generation doublet. The $R_{b}$ bound requires $\lambda_{3}\lesssim 0.04\,M/v\simeq 0.5$ for $M=2$ TeV, comfortably within the $\mathcal{O}(1)$ range. For $\theta_{L}^{b}\sim 0.04$ the new-physics contributions are:
[…]
IX Collider Phenomenology
The common VLQ mass $M$ is the one free parameter of the chain framework that is not fixed by the lattice exponent structure. Because the entry coupling scales as $v\,\epsilon^{A_{i}/9}$ (from the Higgs VEV), the exit coupling scales as $M\,\epsilon^{B_{j}/9}$ (from the VLQ mass term), and the chain propagator scales as $(\kappa_{1}\kappa_{2}\kappa_{3}/M)\,\epsilon^{\Sigma/9}$ [Eq. (18)], the factor of $M$ cancels exactly in the effective Yukawa matrix. All quark mass ratios, mixing angles, and CP phases are therefore $M$-independent—they are determined entirely by the lattice exponents and the $\mathcal{O}(1)$ coefficients.
The chain mass enters the phenomenology only through the SM–VLQ mixing angles $\theta_{L}^{i}\sim\lambda_{i}\,v/M$ [Eq. (57)], which control precision observables, FCNC amplitudes, and collider production rates. From below, current ATLAS and CMS pair-production searches exclude $M\lesssim 1.6$ TeV [27, 28]. From above, there is no sharp theoretical bound: higher $M$ simply reduces $\theta_{L}^{i}$ and decouples the VLQ sector from low-energy observables, while the flavor predictions remain intact. However, for $M\gg v$ the endpoint couplings $\lambda_{i}$ must be correspondingly larger to maintain $\theta_{L}^{b}$ at a level consistent with the measured $Zb\bar{b}$ vertex (Sec. VIII.4); the $R_{b}$ constraint $\theta_{L}^{b}\lesssim 0.04$ requires $\lambda_{3}\lesssim 0.04\,M/v$, which remains comfortably $\mathcal{O}(1)$ for $M\lesssim 5$ TeV but begins to require $\lambda_{3}>1$ above that scale. The benchmark range $M\sim 2$33 TeV adopted throughout this paper therefore represents the window in which the framework is consistent with all current constraints, the endpoint couplings are naturally $\mathcal{O}(1)$, and the VLQ states are accessible at the HL-LHC.
IX.1 Chain mass spectrum
After the flavon develops its VEV, the four chain states $D_{1},\dots,D_{4}$ mix through the nearest-neighbor mass terms. The physical masses are the eigenvalues of $\mathcal{M}_{D}\mathcal{M}_{D}^{\dagger}$ [Eq. (16)]. For degenerate bare masses $M_{a}=M$, the mass eigenvalues of $\mathcal{M}_{D}\mathcal{M}_{D}^{\dagger}$ can be computed exactly. Since $\kappa_{a}\epsilon^{h_{a}/9}$ is $\mathcal{O}(1)$ for the smallest hops (see Sec. IV.5), the splittings are not parametrically small: exact diagonalization gives physical masses ranging from roughly 0.5M0.5\,M to 1.6M1.6\,M for $\kappa_{a}=1$. For $M=2$ TeV the chain states thus span a range of $\sim\!1$33 TeV. The lightest state ($D_{4}$, coupled most strongly to the third generation) is produced first at colliders; heavier chain states may decay via $D_{a}\to D_{a+1}+\text{soft}$ if kinematically accessible, producing cascade signatures with multiple $b$-jets and $W/Z/H$ bosons.
IX.2 Pair production
At the 14 TeV LHC, VLQ pair production proceeds via QCD and is essentially model-independent. Approximate cross sections are $\sim 10$, $1$, $0.1$, and $0.01$ fb for $M_{\rm VLQ}=1.5$, $2.0$, $2.5$, and $3.0$ TeV, respectively [13].
Current ATLAS and CMS searches exclude VLQs with masses below 1.51.51.81.8 TeV (depending on the assumed branching ratios), with the strongest limits reaching $M>1.6$ TeV for singlet $B$-quarks decaying predominantly via $bZ$ and $bH$ [27, 28]. The chain framework with $M\sim 2$ TeV is therefore consistent with all current limits, while remaining within reach of the HL-LHC. With $3~\text{ab}^{-1}$ at the HL-LHC, the discovery reach extends to $M_{\rm VLQ}\approx 2.0$$2.5$ TeV (Fig. 3), depending on decay channels and systematics [13, 27]. At a future 100 TeV $pp$ collider (FCC-hh), QCD pair production would probe VLQ masses up to $\sim\!10$ TeV [29], covering the entire natural parameter space of the chain framework.
IX.3 Single production
Single VLQ production scales as
$\sigma_{\rm single}\propto(\theta_{L}^{b})^{2}\,g_{W}^{2},$ (72)
and is kinematically accessible to higher masses than pair production. For $\theta_{L}^{b}\sim 0.1$ the single-production cross section can exceed pair production above $\sim 2$ TeV, making this a sensitive probe of the chain coupling structure.
IX.4 Decay signatures
The lightest chain state $D_{4}$ (which couples most strongly to the third generation) decays as
$D_{4}\to bZ,\quad bH,\quad tW^{-},$ (73)
with branching ratios approaching $1:1:2$ in the high-mass limit (the Goldstone equivalence theorem). Searches in the $tW$ and $bZ$ channels provide complementary sensitivity [27].
[…]
Figure 3: Approximate VLQ production cross sections at $\sqrt{s}=14$ TeV for pair production (solid) and single production with $\theta_{L}^{b}\sim 0.1$ (dashed). The horizontal dotted line indicates the approximate HL-LHC discovery threshold.
IX.5 Electroweak precision constraints
VLQs that mix with the third generation contribute to the oblique parameters $S$ and $T$ [30]. For a single $SU(2)$-singlet VLQ of mass $M$ mixed with $b_{L}$ at angle $\theta_{L}^{b}$, the leading contributions are
$\Delta T\simeq\frac{3\,(\theta_{L}^{b})^{2}}{16\pi s_{W}^{2}}\frac{m_{t}^{2}}{M_{W}^{2}}\,\ln\frac{M^{2}}{m_{t}^{2}},$ (74)
$\Delta S\simeq\frac{(\theta_{L}^{b})^{2}}{6\pi}\ln\frac{M^{2}}{m_{t}^{2}}.$ (75)
For $\theta_{L}^{b}\sim 0.1$ and $M=2$ TeV: $\Delta T\simeq 0.01$ and $\Delta S\simeq 0.002$, well within the current electroweak fit constraints ($\Delta T<0.1$, $\Delta S<0.1$ at $95\%$ CL) [23]. The additional chain states $D_{1}$$D_{3}$ have even smaller mixing angles with SM quarks and contribute negligibly.
X Comparison with FN, Clockwork, and Chain Approaches
Table 3 contrasts UF with the standard Froggatt–Nielsen mechanism, clockwork/linear dilaton models, and the recent TeV-chain framework of AFHM [12].
Table 3: Comparison of UF with standard FN, clockwork, and the AFHM chain framework [12].
Feature FN Clockwork AFHM chains UF
Exponents Free Free Free Discrete
Discrete symmetry No No No $\mathbb{Z}_{18}$
Algebraic closure No Partial No Yes
TeV states No Yes Yes Yes
FCNC protection Param. Geom. Chain Lattice
Quantitative fit Coarse Coarse Good Exact
Axion quality No No No Yes
Lepton extension Ad hoc Ad hoc Yes Yes
$B$ parameter No No No Yes
The AFHM chain framework [12] shares the core dynamical mechanism with UF: nearest-neighbor VLQ chains at the TeV scale generate hierarchical Yukawa matrices, and chain locality provides FCNC suppression. However, UF is more complete in several respects:
(i) Discrete exponent quantization. The AFHM framework treats the chain couplings as free continuous parameters, allowing any hierarchy to be generated by suitable tuning. UF, by contrast, enforces a $\mathbb{Z}_{18}$ discrete gauge symmetry that quantizes all Yukawa exponents on a ninths lattice. This eliminates continuous parametric freedom in the exponents and makes the hierarchy a structural consequence of the symmetry, not a choice of parameters.
(ii) Exact $\mathcal{O}(1)$ coefficients. With continuous coupling freedom, the AFHM framework necessarily fits the mass spectrum by adjusting magnitudes. The $B$-lattice textures of Eq. (21) reproduce all six quark masses with $\mathcal{O}(1)$ coefficients that are essentially unity ($c_{f}\simeq 1.00$), with no parametric freedom remaining.
(iii) Axion quality protection. The $\mathbb{Z}_{18}$ symmetry that enforces the flavor lattice simultaneously protects the Peccei–Quinn symmetry against gravitational spoiling up to the required operator dimension [21]. This connection between the flavor hierarchy and the strong CP problem is absent in the AFHM framework, where the chain couplings carry no discrete gauge charge.
(iv) Algebraic closure. The ninths lattice is algebraically closed under the operations that appear in the Yukawa construction—addition of exponents (chain hops), multiplication by the expansion parameter (flavon insertions), and the seesaw-like corrections from diagonalization. This closure is a structural consequence of the $\mathbb{Z}_{9}$ discrete gauge symmetry underlying the specific value $B=75/14$, and has no counterpart in a continuous-parameter framework.
(v) Single-parameter organization. The entire quark and lepton mass spectrum is organized by the single parameter $B=75/14$ ($\epsilon=1/B\simeq 0.187$). The AFHM framework, while economical in its field content, requires independent specification of each chain coupling constant.
XI Lepton Sector Extension
The $B$-lattice framework extends naturally to the lepton sector using the same expansion parameter $\epsilon=1/B$ and the same discrete gauge structure [5, 8]. The construction rests on two ingredients: (i) $B$-lattice power counting, which fixes the charged-lepton Yukawa suppressions and neutrino mass eigenvalue ratios, and (ii) an approximate $\mu$$\tau$ symmetry, which fixes the $\mathcal{O}(1)$ structure of the neutrino $23$ block.
XI.1 Charged-lepton texture
The charged-lepton exponent matrix, expressed on the denominator-$18$ lattice, is [8]
$p^{e}=\frac{1}{18}\begin{pmatrix}87&64&40\\ 64&30&36\\ 40&36&0\end{pmatrix},\qquad(Y_{e})_{ij}=c^{e}_{ij}\,\epsilon^{\,p^{e}_{ij}}.$ (76)
This symmetric texture departs from strict charge additivity in the $(2,3)$ sector, where the off-diagonal suppression $\epsilon^{2}$ is stronger than the additive rule would predict; this additional suppression ensures that the charged-lepton contribution to atmospheric mixing remains a perturbative correction to the neutrino-dominant structure.
The diagonal entries fix the charged-lepton mass hierarchy:
$m_{e}:m_{\mu}:m_{\tau}\sim c_{e}\,\epsilon^{29/6}:c_{\mu}\,\epsilon^{5/3}:1,$ (77)
with $c_{e}\simeq 0.93$ and $c_{\mu}\simeq 0.97$—both essentially unity.
The left-handed diagonalization matrix is
$U_{eL}\simeq\begin{pmatrix}1&\epsilon^{17/9}&\epsilon^{20/9}\\ -\epsilon^{17/9}&1&\epsilon^{2}\\ -\epsilon^{20/9}&-\epsilon^{2}&1\end{pmatrix}$, (78)
so the charged-lepton mixing angles are $\theta_{12}^{e}\sim\epsilon^{17/9}\simeq 0.04$, $\theta_{23}^{e}\sim\epsilon^{2}\simeq 0.03$, $\theta_{13}^{e}\sim\epsilon^{20/9}\simeq 0.03$.
In the chain picture, the charged-lepton sector is realized by an analogous vectorlike lepton chain $E_{a}+\bar{E}_{a}$, with the same hop structure $\{1,2,4\}/9$ as the quark chains.
XI.2 Neutrino mass matrix and $\mu$$\tau$ symmetry
Light neutrino masses are generated through the Weinberg operator [5]. The $B$-lattice hierarchy alone does not determine the neutrino mixing pattern; large atmospheric and solar angles require additional $\mathcal{O}(1)$ structure. The FN charge assignment $p^{\nu}_{L_{1}}\gg p^{\nu}_{L_{2}}=p^{\nu}_{L_{3}}$ produces the exponent texture [8]
$p^{\nu}=\frac{1}{9}\begin{pmatrix}9&9&9\\ 9&0&0\\ 9&0&0\end{pmatrix}$, (79)
which suppresses first-generation couplings by $\mathcal{O}(\epsilon)$ relative to the $23$ block. The $\mathcal{O}(1)$ coefficients in the $23$ block must satisfy approximate $\mu$$\tau$ symmetry [31, 32, 33]: $c^{\nu}_{22}\simeq c^{\nu}_{33}$ and $c^{\nu}_{12}\simeq c^{\nu}_{13}$, together with a near-maximal ratio $c^{\nu}_{23}/c^{\nu}_{22}\simeq-1$. This structure can be enforced by a $Z_{2}^{\mu\tau}$ exchange symmetry or by larger discrete groups such as $A_{4}$ or $S_{4}$ [17, 34].
The resulting leading-order neutrino mass matrix is
$\frac{m_{\nu}}{m_{3}}\;\simeq\;\begin{pmatrix}\mathcal{O}(\epsilon)&\mathcal{O}(\epsilon)&\mathcal{O}(\epsilon)\\ \mathcal{O}(\epsilon)&\;\;\;\tfrac{1}{2}&-\tfrac{1}{2}\\ \mathcal{O}(\epsilon)&-\tfrac{1}{2}&\;\;\;\tfrac{1}{2}\end{pmatrix}$, (80)
whose $23$ block has eigenvalues 0 and $1$ (in units of $m_{3}$), giving maximal atmospheric mixing and a near-tribimaximal solar angle. The neutrino mass hierarchy in normal ordering is $m_{3}:m_{2}:m_{1}\sim 1:\epsilon:\epsilon^{2}$, with $m_{3}\simeq 50$ meV, $m_{2}\simeq 9$ meV, $m_{1}\simeq 2$ meV, and $\sum m_{i}\simeq 61$ meV (compatible with cosmological bounds).
The neutrino diagonalization matrix is near-tribimaximal:
$U_{\nu}\simeq\begin{pmatrix}\frac{2}{\sqrt{6}}&\frac{1}{\sqrt{3}}&\mathcal{O}(\epsilon)\\[4.0pt] -\frac{1}{\sqrt{6}}&\frac{1}{\sqrt{3}}&\frac{1}{\sqrt{2}}\\[4.0pt] -\frac{1}{\sqrt{6}}&\frac{1}{\sqrt{3}}&-\frac{1}{\sqrt{2}}\end{pmatrix}$, (81)
with $\theta_{13}^{\nu}\sim\epsilon$ generated by $\mathcal{O}(\epsilon)$ breaking of the $Z_{2}^{\mu\tau}$.
[…]
The PMNS matrix is $U_{\rm PMNS}=U_{eL}^{\dagger}\,U_{\nu}$. Since the charged-lepton angles are small ($\sim\!\epsilon^{17/9}$$\epsilon^{2}$), the large PMNS angles are neutrino-dominant, with $U_{eL}$ providing perturbative corrections.
The reactor element receives contributions from both sectors [8]:
$U_{e3}\simeq s_{13}^{\nu}e^{-i\delta_{\nu}}-\theta_{12}^{e}\,s_{23}^{\nu}\,e^{i\phi_{12}^{e}}-\theta_{13}^{e}\,c_{23}^{\nu}\,e^{i\phi_{13}^{e}}.$ (82)
[…]
$\boxed{\delta\simeq\delta_{\nu}-\arg\!\left(1-r\,e^{-i\Phi_{e}}\right),}$ (83)
and the atmospheric angle receives its dominant correction from the charged-lepton $23$ rotation:
$\boxed{\theta_{23}\simeq\theta_{23}^{\nu}-\theta_{23}^{e}\cos\phi_{23}^{e}.}$ (84)
The single-flavon origin of the $B$-lattice textures ensures [8] that $\phi_{12}^{e}\approx\phi_{23}^{e}$ to within $\sim\!\phi_{0}/9$, where $\phi_{0}=\arg\langle\Phi\rangle$ is the flavon phase. Under this alignment, the sign of the octant shift $(\theta_{23}-45^{\circ})$ correlates directly with the direction of the $\delta$ shift, producing the two-branch prediction.
XI.4 Two-branch prediction
The interplay between Eqs. (83) and (84) yields two normal-ordering solutions [8]:
Table 4: Two-branch prediction for the atmospheric octant and Dirac CP phase. The Jarlskog invariant and effective Majorana mass are branch-independent.
Observable NO-I NO-II Discrim.?
$\theta_{23}$ $43^{\circ}$ $46^{\circ}$ $\checkmark$
$\delta$ $286^{\circ}$ $304^{\circ}$ $\checkmark$
$J_{\rm CP}$ 0.027-0.027 0.025-0.025 $\times$
$m_{\beta\beta}$ [meV] $3.3$ $3.6$ $\times$
NO-I (lower octant) is favored $\sim\!4{:}1$ by a theoretical prior based on the coefficient scan. The framework makes a falsifiable prediction: if future data establish the upper octant, $\delta$ must lie near $304^{\circ}$; if the lower octant, near $286^{\circ}$. Discovery of $\delta$ in the first or second quadrant ($0^{\circ}$$180^{\circ}$) in either octant would rule out the single-$B$ lattice textures with aligned phases. The decisive measurements are $\nu_{\mu}\to\nu_{e}$ appearance rates at DUNE and Hyper-Kamiokande, combined with octant determination from atmospheric neutrinos at IceCube and independent mass-ordering resolution by JUNO.
Three observables are nearly branch-independent and cannot discriminate: $J_{\rm CP}\simeq-0.027$, $m_{\beta\beta}\simeq 3$$4~\text{meV}$, and $\sin^{2}\theta_{13}\simeq 0.022$.
XI.5 Neutrinoless double beta decay
In normal ordering with $m_{3}:m_{2}:m_{1}\sim 1:\epsilon:\epsilon^{2}$, the effective Majorana mass scales as
$m_{\beta\beta}\sim\epsilon\,m_{3}\sim\frac{m_{3}}{B}\simeq 9~\text{meV},$ (85)
within the projected reach of next-generation $0\nu\beta\beta$ experiments. The coefficient scan gives a tighter range $m_{\beta\beta}\simeq 3$44 meV for vanishing Majorana phases (Table 4).
XII Quark–Lepton Connections
The $B$-lattice framework reveals a set of quark–lepton connections that go beyond the usual unification paradigm. In this section we show that the charged-lepton geometric mean ratio (GMR) is directly determined by the lattice parameter $\epsilon$, and that this identification provides a single organizing principle for both CKM and PMNS mixing magnitudes.
XII.1 The charged-lepton GMR and $\epsilon^{3/2}$
Define the charged-lepton geometric mean ratio
$\beta\;\equiv\;\frac{\sqrt{m_{e}^{2}\,m_{\tau}^{2}}}{m_{\mu}^{2}}\;=\;\frac{m_{e}\,m_{\tau}}{m_{\mu}^{2}}.$ (86)
Using $\overline{\rm MS}$ masses at $M_{Z}$, one finds $\beta=0.0806\pm 0.0001$. In the $B$-lattice, the charged-lepton mass scaling $m_{e}:m_{\mu}:m_{\tau}\sim\epsilon^{p_{e}}:\epsilon^{p_{\mu}}:1$ gives
$\beta=\epsilon^{p_{e}-2p_{\mu}}=\epsilon^{29/6-2\times 5/3}=\epsilon^{29/6-10/3}=\epsilon^{3/2},$ (87)
and numerically $\epsilon^{3/2}\simeq 0.0806$—an exact match. The exponent $3/2=Q(L_{1})+Q(L_{2})=1+\tfrac{1}{2}$ is the sum of the first- and second-generation lepton-doublet charges, so $\beta$ is determined by the $Z_{18}$ discrete gauge charges.
Equivalently,
$\frac{1}{\beta}=B^{3/2}\simeq 12.40,$ (88)
providing a direct algebraic link between the charged-lepton mass spectrum and the single $B$-lattice organizing parameter.
Geometric mean mass ratios arise as the natural organizing variables whenever the mass spectrum is a power law of a single parameter: if $m_{i}\propto\epsilon^{p_{i}}$ then $\sqrt{m_{i}m_{j}}/m_{k}=\epsilon^{(p_{i}+p_{j}-2p_{k})/2}$ is itself a power of $\epsilon$, so all such ratios lie on the lattice. This structure is present in any Froggatt–Nielsen-type model with a single flavon [35] and is realized concretely in the UF framework through the multi-messenger sum (Sec. V) that generates the $\mathcal{O}(1)$ coefficients $C^{f}_{ij}$.
XII.2 CKM magnitudes as powers of $\beta$
The $B$-lattice CKM scaling can be reexpressed entirely in terms of $\beta$. Since $\epsilon=\beta^{2/3}$:
$|V_{us}|$ $\sim\epsilon^{8/9}=\beta^{16/27},$ (89)
$|V_{cb}|$ $\sim\epsilon^{17/9}=\beta^{34/27},$ (90)
$|V_{ub}|$ $\sim\epsilon^{10/3}=\beta^{20/9}.$ (91)
The CKM hierarchy is thus controlled by a single number $\beta$ that is itself a ratio of charged-lepton masses. Two cross-checks are immediate:
$\frac{|V_{cb}|}{|V_{us}|}$ $=\beta^{34/27-16/27}=\beta^{2/3},$ (92)
$\frac{|V_{ub}|}{|V_{us}|}$ $=\beta^{20/9-16/27}=\beta^{44/27}.$ (93)
XII.3 PMNS magnitudes as powers of $\beta$
The PMNS mixing angles also organize as powers of $\beta$. A Froggatt–Nielsen-type parameterization with denominator $25$ exponents [8] gives
$\sin\theta_{12}$ $\simeq\beta^{6/25}\simeq 0.547,$ (94)
$\sin\theta_{13}$ $\simeq\beta^{19/25}\simeq 0.148,$ (95)
$\sin\theta_{23}$ $\simeq\beta^{4/25}\simeq 0.668,$ (96)
corresponding to $\theta_{12}\simeq 33^{\circ}$, $\theta_{13}\simeq 8.5^{\circ}$, $\theta_{23}\simeq 42^{\circ}$ (lower octant). The Dirac CP phase is
$\delta\simeq(1/\beta)^{14/25}\simeq 235^{\circ}.$ (97)
These values are consistent with the NO-I solution of the lepton lattice paper (Sec. XI.4).
Three internal relations follow:
  • $\theta_{12}^{\rm PMNS}+\theta_{12}^{\rm CKM}\simeq 33^{\circ}+13^{\circ}\simeq 46^{\circ}$ (QLC) and
    $\theta_{23}^{\rm PMNS}+\theta_{23}^{\rm CKM}\simeq 42^{\circ}+2.4^{\circ}\simeq 44^{\circ}$ (QLC),
    both close to $45^{\circ}$;
  • $\sin\theta_{12}\times\sin\theta_{13}=\beta^{6/25+19/25}=\beta$ (QLC);
  • $\sin\theta_{23}=(\sin\theta_{12})^{2/3}$, since $4/25=(2/3)\times 6/25$.
Here QLC denotes quark–lepton complementarity [36, 37].
The full PMNS magnitude matrix has the FN representation
$|U_{\rm PMNS}|\;\approx\;\begin{pmatrix}\beta^{2/25}&\beta^{6/25}&\beta^{19/25}\\[3.0pt] \beta^{10/25}&\beta^{4/25}&\beta^{4/25}\\[3.0pt] \beta^{8/25}&\beta^{13/50}&\beta^{3/25}\end{pmatrix}$. (98)
Since $\beta=\epsilon^{3/2}$, each exponent $p/25$ translates to an $\epsilon$-exponent $3p/50$, connecting the PMNS structure back to the $B$-lattice.
Thus both CKM and PMNS mixing hierarchies are controlled by rational powers of the same charged-lepton GMR $\beta=\epsilon^{3/2}$.
Relation to the near-tribimaximal structure.
The $\beta$-power parameterization and the near-tribimaximal (TBM) construction of Sec. XI are complementary descriptions of the same physics. Exact TBM predicts $\sin\theta_{12}=1/\sqrt{3}\simeq 0.577$, $\sin\theta_{13}=0$, and $\sin\theta_{23}=1/\sqrt{2}\simeq 0.707$. The $\beta$-power values $\sin\theta_{12}\simeq 0.547$, $\sin\theta_{13}\simeq 0.148$, $\sin\theta_{23}\simeq 0.668$ depart from TBM by $\sim\!5\%$—precisely the $\mathcal{O}(\epsilon)$ magnitude of the charged-lepton corrections $\theta^{e}_{ij}\sim\epsilon^{17/9}$$\epsilon^{2}$ derived in Sec. XI. Thus the near-TBM framework provides the structural origin of the PMNS pattern (approximate $\mu$$\tau$ symmetry in $U_{\nu}$, broken by $U_{eL}$), while the $\beta$-power representation captures the algebraic result after the charged-lepton corrections have acted. Numerically, the two parameterizations agree at the sub-percent level on all three mixing angles:
TBM Near-TBM $\beta^{p}$ NuFIT
(exact) (NO-I) (NO)
$\sin\theta_{12}$ 0.577 0.545 0.547 0.550
$\sin\theta_{13}$ 0 0.148 0.148 0.148
$\sin\theta_{23}$ 0.707 0.682 0.668 0.672
$\theta_{23}$ octant max lower lower lower
Both select the NO-I (lower-octant) solution and are consistent with NuFIT 6.0.
XII.4 Dual mass representations of CKM elements
Because every quark mass is a definite power of $\epsilon=1/B$, each CKM magnitude can be expressed equivalently in terms of down-type mass ratios ($d,s,b$), up-type mass ratios ($u,c,t$), or mixed combinations. We highlight several notable consequences; a systematic treatment is given elsewhere.
Dual representations.
The Cabibbo element admits four equivalent forms:
$|V_{us}|=\Bigl(\frac{d}{s}\Bigr)^{\!1/2}=\frac{c}{b}=\Bigl(\frac{u}{t}\Bigr)^{\!1/8}=\Bigl[\frac{(ut)^{1/2}}{c}\Bigr]^{4}.$ (99)
Each form equals $\epsilon^{8/9}\simeq 0.225$. Similarly,
$|V_{cb}|=\Bigl(\frac{u}{c}\Bigr)^{\!1/2}=\frac{db}{ut},\qquad|V_{ub}|=\frac{c}{t}=\frac{ds}{uc}.$ (100)
Mass sum rules.
Equating dual representations yields algebraic relations among quark masses. Two illustrative ones are
$uc^{2}=dsb,\qquad\frac{(ub)^{1/2}}{s}=1,$ (101)
both satisfied to sub-permille accuracy by the $B$-lattice benchmark masses. The first can be written as
$\frac{tcu}{bsd}=\frac{1}{V_{ub}}=B^{10/3}\simeq 269,$ (102)
expressing the full quark mass product ratio as a power of $B$. Also note that
$B=\left(\frac{tc}{sd}\right)^{\!1/8}.$ (103)
These identities are not numerological coincidences: they are algebraic consequences of the fact that all six quark masses are determined by a single parameter $B$ and the discrete $Z_{9}$ charges. In the $B$-lattice, the masses scale as $(u,c,t)=t\,(\epsilon^{64/9},\,\epsilon^{10/3},\,1)$ and $(d,s,b)=b\,(\epsilon^{37/9},\,\epsilon^{7/3},\,1)$, so any ratio of mass monomials reduces to a rational power of $\epsilon$. A complete catalog of the resulting dual CKM representations and geometric mean relations will be presented in a forthcoming companion paper.
XII.5 Color-factor mass relation and the golden ratio
Two additional quark–lepton mass relations follow from the $B$-lattice charge assignments:
  1. 1.
    Color factor. The geometric means of the down-quark and charged-lepton mass-squares satisfy
    $\bigl(d^{2}s^{2}b^{2}\bigr)^{1/3}=3\,\bigl(e^{2}\mu^{2}\tau^{2}\bigr)^{1/3},$ (104)
    to sub-percent accuracy. The factor of 3 suggests an underlying SU(3)c color-weight connection between the two sectors.
  2. 2.
    Golden ratio. The $b$-quark to $\tau$-lepton mass ratio is
    $\frac{m_{b}}{m_{\tau}}=\frac{m_{b}^{\rm ref}}{m_{\tau}^{\overline{\rm MS}}(M_{Z})}\simeq\varphi\equiv\frac{1+\sqrt{5}}{2}\simeq 1.618,$ (105)
    as previously noted in Paper I [6]. This relation is used to set the reference scale $m_{b}^{\rm ref}=\varphi\,m_{\tau}^{\overline{\rm MS}}(M_{Z})$ in the quark-mass benchmark of Sec. V.
These relations, together with the $\beta$ parameterization of CKM and PMNS magnitudes, provide evidence that the quark and lepton mass spectra are organized by a common principle rooted in the $B$-lattice charge assignments.
XIII Discrete Anomaly Cancellation
A discrete gauge symmetry is physical—and in particular can protect the axion quality—only if it is free of gauge anomalies. For a $Z_{N}$ symmetry, the Ibáñez–Ross conditions [4] are the discrete analogs of continuous anomaly cancellation. Writing $q_{i}$ for the $Z_{N}$ charge of each left-handed Weyl fermion, the conditions require
$A_{3}$ $\equiv\sum_{\rm color}\!q_{i}\,d_{2}(i)\equiv 0\pmod{N}$, (106)
$A_{2}$ $\equiv\sum_{\rm weak}\!q_{i}\,d_{3}(i)\equiv 0\pmod{N}$, (107)
$A_{\rm grav}$ $\equiv\sum_{\rm all}\!q_{i}\,d_{3}(i)\,d_{2}(i)\equiv 0\pmod{N}$, (108)
where $d_{2}$ and $d_{3}$ denote the SU(2)L and SU(3)c representation dimensions, and the sums run over the indicated subsets of fermions.
XIII.1 $Z_{18}$ charge assignments
The complete $Z_{18}$ charge table follows from the $B$-lattice FN charges determined in Secs. V and XI [5, 21], expressed as integer charges $q_{18}\equiv 18\,Q$:
[…]
Here $N_{i}$ are right-handed (sterile) neutrinos and $H$ is the SM Higgs doublet. The third generation of all SM fields carries zero charge, so the anomaly sums receive contributions only from the first two generations. The $Z_{9}$ subgroup (denominator-$9$ lattice) controls quark-sector exponents, while the full $Z_{18}$ is required to accommodate the half-integer lepton-doublet charge $Q(L_{2})=\tfrac{1}{2}$.
Uniqueness of the charge table.
It is important to emphasize that the charges in Eq. (109) are not one of many possible anomaly-free solutions selected to match the data. The logical sequence is reversed: the quark-sector charges $Q(Q_{i})$ and $Q(f^{c}_{j})$ are uniquely fixed by the observed mass ratios and the value $B=75/14$ (Sec. V), the lepton-sector charges are similarly determined by the charged-lepton and neutrino mass hierarchies (Sec. XI), and the anomaly conditions are then a nontrivial consistency check on these data-determined charges. Because the charges are overdetermined by the mass spectrum, one cannot freely adjust them to satisfy the anomaly conditions—the fact that they do so is a structural property of the $B$-lattice, not a consequence of parameter tuning.
Bottom-up origin and uniqueness. The charges in Eq. (109) are not free parameters adjusted to satisfy anomaly cancellation. They are determined bottom-up by the observed fermion mass spectrum through the $B$-lattice formula $m_{f}=m_{b}^{\rm ref}\cdot\epsilon^{n_{f}}$ (Paper I [6]): the diagonal exponents $n_{f}$ are read off from the measured quark and lepton mass ratios, and the factorization $p^{f}_{ij}=Q(Q_{i})+Q(f^{c}_{j})$ then determines all charges up to the convention $Q(\cdot_{3})=0$ for the third generation. Within the $B$-lattice framework, these charges are therefore unique—they are the only set that simultaneously reproduces the six quark mass ratios, three CKM magnitudes, and three charged-lepton masses with $\mathcal{O}(1)$ coefficients $c_{f}\simeq 1.00$ and a single expansion parameter $\epsilon=1/B$. That the resulting $q_{18}$ values happen to satisfy the Ibáñez–Ross anomaly conditions (106)–(108) (modulo the Green–Schwarz shift of Sec. XIII.3) is a nontrivial consistency check—a prediction of the framework, not an input.
XIII.2 Neutrino masses from the charge table
The $Z_{18}$ charges determine the neutrino mass texture through two complementary mechanisms.
Weinberg operator.
The dimension-five operator $L_{i}L_{j}HH/\Lambda$ carries $Z_{18}$ charge $Q(L_{i})+Q(L_{j})$, so the required number of flavon insertions gives the exponent matrix $p^{\nu}_{ij}=Q(L_{i})+Q(L_{j})$ of Eq. (79).
Type-I seesaw.
With $Q(N_{i})=0$, the Dirac Yukawa matrix scales as $(Y_{\nu})_{ij}\sim\epsilon^{Q(L_{i})}$ (independent of the right-handed generation index), and the Majorana mass matrix $M_{N}$ is unsuppressed. Integrating out $N$ via the seesaw formula $(m_{\nu})_{ij}\sim v^{2}(Y_{\nu}M_{N}^{-1}Y_{\nu}^{T})_{ij}$ reproduces
$(m_{\nu})_{ij}\sim\frac{v^{2}}{\Lambda_{N}}\,\epsilon^{Q(L_{i})+Q(L_{j})},$ (110)
recovering the Weinberg-operator texture with the identification $\Lambda_{N}\sim\Lambda$. The diagonal entries give the mass eigenvalue hierarchy
$m_{1}:m_{2}:m_{3}\sim\epsilon^{2Q(L_{1})}:\epsilon^{2Q(L_{2})}:\epsilon^{2Q(L_{3})}=\epsilon^{2}:\epsilon:1,$ (111)
in agreement with normal ordering.
The vanishing $Z_{18}$ charge of $N_{i}$ has two further consequences: the right-handed neutrinos are SM singlets with zero discrete charge, so they do not contribute to any of the anomaly sums (106)–(108); and their Majorana mass $M_{N}\sim\Lambda$ is unsuppressed by the flavon, placing the seesaw scale at $\Lambda\sim 5\times 10^{11}$ GeV, consistent with the axion decay constant $f_{a}$ discussed in Sec. XIV.
XIII.3 Green–Schwarz cancellation and the axion
The SM fermion charges above do not by themselves satisfy Eqs. (106)–(108); the residual anomalies $A_{a}\not\equiv 0\pmod{18}$ are nonzero. This is a generic feature of FN models and is resolved by the Green–Schwarz (GS) mechanism: the $Z_{18}$ discrete gauge symmetry descends from a continuous $U(1)_{\rm FN}$ that is spontaneously broken by the flavon VEV $\langle\Phi\rangle$. The $U(1)_{\rm FN}$ mixed anomalies are cancelled by a GS axion, and this cancellation persists in the discrete $Z_{N}$ remnant [4, 38].
The connection to the strong CP problem is immediate: the same PQ axion that solves the strong CP problem (Sec. XIV) can serve as the GS field that cancels the discrete gauge anomalies. This identification is natural in the “Flavor in Ninths” framework [21], where the $Z_{18}$ discrete gauge symmetry simultaneously
  1. 1.
    enforces the ninths-quantized lattice exponents,
  2. 2.
    protects the PQ symmetry from Planck-suppressed operators, and
  3. 3.
    is rendered anomaly-free by the PQ/GS axion.
The VLQ chain fields ($D_{a}+\bar{D}_{a}$, $U_{a}+\bar{U}_{a}$) are vectorlike under all gauge groups including $Z_{18}$ and therefore do not contribute to the anomaly sums.
XIII.4 Spectator fermion alternative
If the GS mechanism is not invoked, anomaly cancellation can alternatively be achieved by introducing spectator fermions at a high scale $\Lambda\gg M_{\rm VLQ}$ that carry appropriate $Z_{18}$ charges but are SM singlets. Such spectators are common in UV completions of FN models and decouple from low-energy physics. In either case, the anomaly-free $Z_{18}$ is a legitimate discrete gauge symmetry, and its protection of the axion quality and the lattice exponent structure is on firm theoretical footing.
XIV Connection to the Axion Quality Problem
The “Flavor in Ninths” paper [21] demonstrated that the same discrete $Z_{N}$ symmetry enforcing the ninths lattice can simultaneously address the Peccei–Quinn (PQ) axion quality problem [39]. The axion quality problem—the requirement that Planck-suppressed operators violating the global $U(1)_{\rm PQ}$ symmetry be forbidden to sufficiently high dimension—was identified by Kamionkowski and March-Russell [40], Holman et al. [41], and Barr and Seckel [42], and has been studied quantitatively in the context of the string landscape by Baer, Barger, and Sengupta [43] and in gauge-protected axion models by Bhattiprolu and Martin [44]. In the UF construction, the FN flavon $\Phi$ is identified with the PQ scalar, and the $Z_{18}$ discrete gauge symmetry forbids all PQ-violating Planck-suppressed operators below dimension $9$.
The idea of connecting the flavor symmetry to axion quality has been pursued independently by Babu, Chandrasekar, and Tavartkiladze [20], who use a gauged continuous $U(1)_{F}$ symmetry to realize the FN mechanism while an accidental global $U(1)_{\rm PQ}$ emerges as a byproduct, yielding a high-quality Dine–Fischler–Srednicki–Zhitnitsky (DFSZ)-type axion [45, 46]. UF achieves the same connection through a discrete $\mathbb{Z}_{18}$ gauge symmetry rather than a continuous $U(1)_{F}$. The discrete realization has two structural advantages: it fixes the Yukawa exponents to the ninths lattice (rather than allowing continuous charge freedom), and it avoids the need for the continuous gauge boson $Z^{\prime}_{F}$ and its associated phenomenological constraints.
This connection carries over to the chain framework: the $Z_{9}$ (or $Z_{18}$) symmetry that enforces nearest-neighbor locality along the VLQ chain is the same symmetry that protects the axion potential. The chain construction therefore inherits the axion-quality protection automatically, without requiring any additional symmetry.
In this picture, the axion decay constant is set by the flavon VEV,
$f_{a}=\langle\Phi\rangle=\epsilon\cdot\Lambda,$ (112)
and the UV cutoff $\Lambda$ is identified with $\Lambda\sim f_{a}/\epsilon\sim 5\times 10^{11}$ GeV for $f_{a}\sim 10^{11}$ GeV. The VLQ chain mass scale $M\sim\mathcal{O}(\text{TeV})$ is then a separate, lower scale, and the chain states are much lighter than the flavon/messenger sector.
The axion in this mass window ($f_{a}\sim 10^{11}$ GeV) is a viable dark matter candidate. The continued absence of weakly interacting massive particle (WIMP) signals in direct detection experiments [47, 48, 49] suggests that WIMPs—if they were ever thermally produced—may have decayed away via small symmetry-violating couplings, leaving axions as the dominant component of dark matter today [50, 51]. In the UF framework this outcome is natural: the same $Z_{18}$ discrete gauge symmetry that protects the axion quality also forbids the lowest-dimension operators that would destabilize the lightest supersymmetric particle in SUSY extensions, providing a unified explanation for both the flavor hierarchy and the dark matter composition.
The triple unification of flavor hierarchy, axion quality, and chain locality through a single discrete gauge symmetry constitutes one of the most distinctive features of UF.
XV Renormalization Group Stability
The $B$-lattice textures are defined at the chain scale $\mu\sim M$. Renormalization group (RG) running from $M$ down to $M_{Z}$ modifies the Yukawa couplings through gauge and Yukawa loop corrections. Two key features ensure stability of the lattice structure:
First, the dominant RG effect—QCD running—is flavor universal: it rescales all down-type Yukawa entries by a common factor $\eta_{\rm QCD}\simeq({\alpha_{s}(M)}/{\alpha_{s}(M_{Z})})^{4/\beta_{0}}$, where $\beta_{0}=(33-2n_{f})/3$ is the one-loop QCD beta-function coefficient with $n_{f}$ active quark flavors. This shifts the overall mass scale (absorbed into $m_{b}$) but preserves all mass ratios and mixing angles.
Second, the Yukawa self-corrections are hierarchical: the top Yukawa dominates, modifying primarily $(Y_{u})_{33}$, while the down-sector texture receives corrections of order $y_{t}^{2}/(16\pi^{2})\cdot\ln(M/M_{Z})\sim 0.02$ per entry. Since the lattice exponent differences are $\mathcal{O}(1)$ in units of $\epsilon$, the RG corrections are parametrically smaller ($\sim y_{t}^{2}/16\pi^{2}\ll\epsilon$) and do not disturb the power-counting structure.
Thus the chain-generated textures at $\mu=M$ flow smoothly to the $B$-lattice textures at $\mu=M_{Z}$ identified in Papers I–IV, with corrections that are absorbed into the $\mathcal{O}(1)$ coefficients.
XVI Discussion
Papers I–IV of the $B$-lattice program established the lattice exponent structure from the observed fermion mass and mixing hierarchies. The present work demonstrates that this structure is not merely a convenient parametrization, but emerges dynamically from a well-defined UV completion:
  1. 1.
    The $Z_{N}$ discrete gauge symmetry quantizes the FN exponents to the ninths lattice $\mathcal{L}$.
  2. 2.
    The $Z_{2}^{\rm(NN)}$ locality symmetry restricts the allowed operator products to nearest-neighbor chain compositions.
  3. 3.
    The Path-Sum Theorem converts these restrictions into a definite prediction: Yukawa exponents are sums of discrete hop contributions.
  4. 4.
    The TeV-scale VLQ masses provide the physical cutoff $\Lambda\sim M$, making the framework testable at colliders.
  5. 5.
    The discrete anomaly of $Z_{18}$ is cancelled by the Green–Schwarz mechanism, identifying the GS axion with the PQ axion that solves the strong CP problem. The specific group order $N=18$ is not a free choice: it is the minimal $Z_{N}$ that simultaneously accommodates the half-integer lepton doublet charge $Q(L_{2})=\tfrac{1}{2}$, the ninths-quantized quark exponents, and the axion quality requirement that PQ-violating operators be forbidden below dimension $N/2=9$.
The lattice hierarchy is therefore structural rather than merely parametric: it is a consequence of the symmetry and field content, not a choice of parameters. The fact that the same discrete gauge group that fixes the flavor exponents also renders the axion potential gravity-safe elevates $\mathbb{Z}_{18}$ from a bookkeeping device to a symmetry with cosmological consequences. The key predictions and their experimental tests are collected in Table 5.
Table 5: Key predictions and experimental tests of the UF chain framework.
Observable Prediction Test
$m_{d}:m_{s}:m_{b}$ $\epsilon^{37/9}:\epsilon^{7/3}:1$ PDG masses
$m_{u}:m_{c}:m_{t}$ $\epsilon^{64/9}:\epsilon^{10/3}:1$ PDG masses
$|V_{us}|:|V_{cb}|:|V_{ub}|$ $\epsilon^{8/9}:\epsilon^{17/9}:\epsilon^{10/3}$ PDG CKM
$J_{\rm CKM}$ $\sim 3\times 10^{-5}$ PDG
$M_{\rm VLQ}$ 2233 TeVa HL-LHC
$D_{4}$ decays $bZ:bH:tW\sim 1:1:2$ HL-LHC
Chain cascade multi-$b$ + $W/Z/H$ HL-LHC
$\Delta T$, $\Delta S$ $\lesssim 0.01$ EW fit
$\delta R_{b}/R_{b}$ $\lesssim 0.4\%$ LEP $R_{b}$
$|\delta C_{7}/C_{7}^{\rm SM}|$ $\sim 10^{-3}$ $b\!\to\!s\gamma$
$C_{10}^{\rm NP}$ $\sim-0.3$ $B_{s}\!\to\!\mu^{+}\mu^{-}$
$C_{K\bar{K}}$ $\sim 10^{-20}~\text{GeV}^{-2}$ Kaon mixing
Axion quality $Z_{18}$ protected Haloscopes
$m_{e}:m_{\mu}:m_{\tau}$ $\epsilon^{13/3}:\epsilon^{2/3}:1$ PDG masses
$m_{3}:m_{2}:m_{1}$ $1:\epsilon:\epsilon^{2}$ Oscillations
[…]
$m_{\beta\beta}$ $\sim 9$ meV $0\nu\beta\beta$
aBenchmark range (Sec. IX), not a lattice prediction.
XVII Conclusions
Unified Flavor provides the dynamical ultraviolet completion of the $B$-lattice hierarchy program. A single flavon $\Phi$ with $\epsilon=1/B=14/75$ enforces discrete ninths-quantized Yukawa exponents, while TeV-scale nearest-neighbor VLQ chains generate the effective Yukawa couplings as algebraic path sums. Each effective Yukawa amplitude factorizes into entry, chain-propagation, and exit factors controlled by the discrete $Z_{9}$ charges of the SM fields, with the mass hierarchy encoded in the exponents $\epsilon^{Q(Q_{i})+Q(f^{c}_{j})}$. A multi-messenger structure—coherent contributions from several chain configurations—generates $\mathcal{O}(1)$ complex coefficients whose phases are the physical origin of CKM CP violation.
The resulting textures reproduce both quark mass hierarchies ($m_{d}:m_{s}:m_{b}\sim\epsilon^{37/9}:\epsilon^{7/3}:1$, $m_{u}:m_{c}:m_{t}\sim\epsilon^{64/9}:\epsilon^{10/3}:1$) with $\mathcal{O}(1)$ coefficients $c_{f}\simeq 1.00$ for all six quarks, the CKM mixing hierarchy ($|V_{us}|:\!|V_{cb}|:\!|V_{ub}|\sim\epsilon^{8/9}:\!\epsilon^{17/9}:\!\epsilon^{10/3}$), and the CP-violating phase through adjacent-exponent interference—all from a single parameter $B=75/14$.
The chain locality simultaneously provides a built-in Glashow–Iliopoulos–Maiani-like suppression of FCNCs: tree-level $\Delta F=2$ operators lie orders of magnitude below current bounds, one-loop dipole operators ($b\to s\gamma$) shift $C_{7}$ by less than $1\%$, and the dominant new-physics effect—the tree-level $Z$-FCNC contribution to $B_{s}\to\mu^{+}\mu^{-}$—is consistent with current precision for VLQ masses $M\gtrsim 2$ TeV. The framework requires multi-TeV VLQs decaying as $D\to bZ$, $bH$, $tW^{-}$, accessible at the HL-LHC with $3~\text{ab}^{-1}$. The same $\mathbb{Z}_{18}$ discrete gauge symmetry that enforces the lattice structure simultaneously protects the Peccei–Quinn axion quality, with the discrete anomaly cancelled by the Green–Schwarz mechanism through the identification of the GS axion with the QCD axion. This provides a triple unification of the flavor hierarchy, CP violation, and the strong CP problem in which a single discrete gauge group—whose order is fixed by the joint requirements of the ninths lattice and the lepton charge assignments—determines the Yukawa exponent algebra, forbids dangerous Planck-suppressed PQ-violating operators, and is rendered anomaly-free by the axion itself.
[…]
Unified Flavor completes the $B$-lattice series by demonstrating that the phenomenologically successful lattice exponent algebra emerges from a definite symmetry structure and field content, rendering the fermion mass hierarchy a structural consequence of the UV theory rather than a parametric accident.
Appendix A Step-by-Step Block Inversion of the Chain Mass Matrix
We derive the corner element $(\mathcal{M}^{-1})_{N1}$ for an $N$-site upper-triangular chain mass matrix [Eq. (13)] and show that the result is exact—valid for arbitrary magnitudes of the off-diagonal entries $m_{a,a+1}$. The key observation is that the inverse of a nonsingular upper-triangular matrix is also upper-triangular and is computed exactly by back-substitution [22]. Equivalently, Cramer’s rule gives the $(1,N)$ cofactor as the determinant of an $(N\!-\!1)\times(N\!-\!1)$ lower-triangular submatrix whose diagonal entries are precisely $m_{12},m_{23},\dots,m_{N-1,N}$, yielding the product formula without any series expansion.
A.1 $N=2$
For two sites:
$\mathcal{M}_{2}=\begin{pmatrix}M_{1}&m_{12}\\ 0&M_{2}\end{pmatrix},\qquad\mathcal{M}_{2}^{-1}=\begin{pmatrix}M_{1}^{-1}&-m_{12}/(M_{1}M_{2})\\ 0&M_{2}^{-1}\end{pmatrix}$. (113)
Thus
$(\mathcal{M}_{2}^{-1})_{21}=0,\qquad(\mathcal{M}_{2}^{-1})_{12}=-\frac{m_{12}}{M_{1}M_{2}}.$ (114)
Note: the physically relevant element for the Yukawa is $(\mathcal{M}^{-1})_{N1}$, which connects the last site (coupled to $q_{L}$) to the first site (coupled to $d_{R}$).
A.2 $N=3$
Writing $\mathcal{M}_{3}$ in $2\times 2$ block form and using the Schur complement:
$(\mathcal{M}_{3}^{-1})_{31}=\frac{m_{12}\,m_{23}}{M_{1}\,M_{2}^{2}\,M_{3}}\times M_{2}=\frac{m_{12}\,m_{23}}{M_{1}\,M_{2}\,M_{3}}.$ (115)
A.3 General NN
By induction, the $N$-site corner element is
$(\mathcal{M}^{-1})_{N1}=(-1)^{N-1}\frac{\prod_{k=1}^{N-1}m_{k,k+1}}{\prod_{a=1}^{N}M_{a}}.$ (116)
For degenerate masses $M_{a}=M$ and $m_{k,k+1}=\kappa_{k}M\,\epsilon^{h_{k}/9}$:
$(\mathcal{M}^{-1})_{N1}=\frac{(-1)^{N-1}}{M}\prod_{k=1}^{N-1}\kappa_{k}\,\epsilon^{h_{k}/9}=\frac{(-1)^{N-1}\prod\kappa_{k}}{M}\,\epsilon^{\Sigma/9}.$ (117)
The overall sign is absorbed into the definition of the Yukawa coupling. Importantly, no step in this derivation assumes $m_{k,k+1}\ll M_{k}$; the result is an algebraic identity for any nonsingular upper-triangular matrix. The condition $\kappa_{a}\epsilon^{h_{a}/9}\ll 1$ is therefore not required for the validity of Eqs. (14)–(15) or the Path-Sum Theorem.
Appendix B Numerical Consistency Checks
With $B=75/14$ and $\epsilon=1/B\simeq 0.1867$, Table 6 collects the numerical values of the principal $B$-lattice quantities.
Table 6: Numerical values of $B$-lattice quantities relevant to the chain framework.
Quantity Expression Value
$\epsilon$ $14/75$ $0.1867$
$\epsilon$ $\theta_{12}^{dL}$ scaling $0.187$
$\epsilon^{2}$ $\theta_{23}^{dL}$ scaling $0.035$
$\epsilon^{3}$ $\theta_{13}^{dL}$ scaling $0.0065$
$\epsilon^{1/3}$ $\theta_{23}^{dR}$ scaling $0.571$
$\epsilon^{7/9}$ Chain hop suppression $0.282$
$\epsilon^{8/9}$ $|V_{us}|$ scaling $0.225$
$\epsilon^{7/3}$ $m_{s}/m_{b}$ scaling $0.020$
$\epsilon^{17/9}$ $|V_{cb}|$ scaling $0.042$
$\epsilon^{10/3}$ $m_{c}/m_{t}$, $|V_{ub}|$ scaling $3.7\times 10^{-3}$
$\epsilon^{37/9}$ $m_{d}/m_{b}$ scaling $1.0\times 10^{-3}$
$\epsilon^{64/9}$ $m_{u}/m_{t}$ scaling $6.6\times 10^{-6}$
$\epsilon^{7/9}/M$ Chain suppression $(1.4\times 10^{-4})~\text{GeV}^{-1}$
($M=2$ TeV)
Appendix C Perturbative Diagonalization of the Full $B$-Lattice Textures
We carry out the perturbative diagonalization of the Yukawa matrices $(Y_{f})_{ij}=c^{f}_{ij}\,\epsilon^{\,p^{f}_{ij}}$ using the exponent matrices of Eq. (21), keeping track of the $\mathcal{O}(1)$ coefficients $c^{f}_{ij}$. All results are given at leading order in $\epsilon$.
C.1 Down sector
The down-type Yukawa matrix has the hierarchical structure
$Y_{d}\sim\begin{pmatrix}c_{11}^{d}\,\epsilon^{37/9}&c_{12}^{d}\,\epsilon^{10/3}&c_{13}^{d}\,\epsilon^{3}\\[3.0pt] c_{21}^{d}\,\epsilon^{28/9}&c_{22}^{d}\,\epsilon^{7/3}&c_{23}^{d}\,\epsilon^{2}\\[3.0pt] c_{31}^{d}\,\epsilon^{10/9}&c_{32}^{d}\,\epsilon^{1/3}&c_{33}^{d}\end{pmatrix}$. (118)
Third generation. At zeroth order, the $(3,3)$ element dominates: $y_{b}=|c_{33}^{d}|$.
2–3 rotation. The leading left-handed rotation that removes $Y_{23}$ is
$\theta_{23}^{dL}\approx\frac{|c_{23}^{d}|}{|c_{33}^{d}|}\,\epsilon^{2}.$ (119)
The corresponding right-handed angle is $\theta_{23}^{dR}\approx(|c_{32}^{d}|/|c_{33}^{d}|)\,\epsilon^{1/3}$.
Second generation. After the 2233 rotation, the effective $(2,2)$ element is
$y_{s}\simeq\left|c_{22}^{d}-\frac{c_{23}^{d}c_{32}^{d}}{c_{33}^{d}}\right|\epsilon^{7/3}.$ (120)
Thus $m_{s}/m_{b}\sim\epsilon^{7/3}$ with an $\mathcal{O}(1)$ effective coefficient.
1–2 rotation. The left-handed angle removing the residual $(1,2)$ element is
$\theta_{12}^{dL}\approx\frac{|c_{12}^{d,\rm eff}|}{|c_{22}^{d,\rm eff}|}\,\epsilon^{10/3-7/3}=\frac{|c_{12}^{d,\rm eff}|}{|c_{22}^{d,\rm eff}|}\,\epsilon,$ (121)
where $c^{d,\rm eff}$ denotes the coefficient after the 2233 rotation.
First generation. The lightest eigenvalue is
$y_{d}\simeq|c_{11}^{d,\rm eff}|\,\epsilon^{37/9},$ (122)
giving $m_{d}/m_{b}\sim\epsilon^{37/9}$.
Summary. The down-sector eigenvalues and mixing angles are:
$y_{b}$ $\sim 1,$ $y_{s}$ $\sim\epsilon^{7/3},$ $y_{d}$ $\sim\epsilon^{37/9},$ (123)
$\theta_{23}^{dL}$ $\sim\epsilon^{2},$ $\theta_{12}^{dL}$ $\sim\epsilon,$ $\theta_{13}^{dL}$ $\sim\epsilon^{3},$ (124)
$\theta_{23}^{dR}$ $\sim\epsilon^{1/3},$ $\theta_{12}^{dR}$ $\sim\epsilon^{7/9},$ $\theta_{13}^{dR}$ $\sim\epsilon^{10/9}.$ (125)
The left-handed angles scale as $\epsilon^{|Q(Q_{i})-Q(Q_{j})|}$ and the right-handed angles as $\epsilon^{|Q(d^{c}_{i})-Q(d^{c}_{j})|}$.
C.2 Up sector
The up-type Yukawa matrix is
$Y_{u}\sim\begin{pmatrix}c_{11}^{u}\,\epsilon^{64/9}&c_{12}^{u}\,\epsilon^{13/3}&c_{13}^{u}\,\epsilon^{3}\\[3.0pt] c_{21}^{u}\,\epsilon^{55/9}&c_{22}^{u}\,\epsilon^{10/3}&c_{23}^{u}\,\epsilon^{2}\\[3.0pt] c_{31}^{u}\,\epsilon^{37/9}&c_{32}^{u}\,\epsilon^{4/3}&c_{33}^{u}\end{pmatrix}$. (126)
Proceeding identically:
Third generation. $y_{t}=|c_{33}^{u}|$.
2–3 rotation.
$\theta_{23}^{uL}\approx\frac{|c_{23}^{u}|}{|c_{33}^{u}|}\,\epsilon^{2},\qquad\theta_{23}^{uR}\approx\frac{|c_{32}^{u}|}{|c_{33}^{u}|}\,\epsilon^{4/3}.$ (127)
Second generation.
$y_{c}\simeq\left|c_{22}^{u}-\frac{c_{23}^{u}c_{32}^{u}}{c_{33}^{u}}\right|\epsilon^{10/3}.$ (128)
1–2 rotation.
$\theta_{12}^{uL}\approx\frac{|c_{12}^{u,\rm eff}|}{|c_{22}^{u,\rm eff}|}\,\epsilon^{13/3-10/3}=\frac{|c_{12}^{u,\rm eff}|}{|c_{22}^{u,\rm eff}|}\,\epsilon.$ (129)
First generation. $y_{u}\simeq|c_{11}^{u,\rm eff}|\,\epsilon^{64/9}$.
Summary.
$y_{t}$ $\sim 1,$ $y_{c}$ $\sim\epsilon^{10/3},$ $y_{u}$ $\sim\epsilon^{64/9},$ (130)
$\theta_{23}^{uL}$ $\sim\epsilon^{2},$ $\theta_{12}^{uL}$ $\sim\epsilon,$ $\theta_{13}^{uL}$ $\sim\epsilon^{3},$ (131)
$\theta_{23}^{uR}$ $\sim\epsilon^{4/3},$ $\theta_{12}^{uR}$ $\sim\epsilon^{25/9},$ $\theta_{13}^{uR}$ $\sim\epsilon^{37/9}.$ (132)
The right-handed angles scale as $\epsilon^{|Q(u^{c}_{i})-Q(u^{c}_{j})|}$ and are substantially more hierarchical than in the down sector, reflecting the wider spacing of the up-type right-handed charges.
C.3 Relation to the Hermitian-square approach
An alternative route to the left-handed mixing angles and mass eigenvalues proceeds through the Hermitian square $H_{f}\equiv Y_{f}Y_{f}^{\dagger}$, whose eigenvalues are $y_{f_{i}}^{2}$ and whose diagonalizing unitary is $U_{fL}$. For the factorized texture $(Y_{f})_{ij}=c^{f}_{ij}\,\epsilon^{Q(Q_{i})+Q(f^{c}_{j})}$ the $(i,j)$ entry of $H_{f}$ is
$(H_{f})_{ij}=\sum_{k=1}^{3}c^{f}_{ik}\,(c^{f}_{jk})^{*}\,\epsilon^{\,Q(Q_{i})+Q(Q_{j})+2Q(f^{c}_{k})}.$ (133)
Because $Q(f^{c}_{3})=0$, the $k=3$ term dominates each entry, giving
$(H_{f})_{ij}\simeq c^{f}_{i3}\,(c^{f}_{j3})^{*}\,\epsilon^{\,Q(Q_{i})+Q(Q_{j})}+\mathcal{O}(\epsilon^{2Q(f^{c}_{2})}).$ (134)
This leading piece is rank one: it is the outer product of the third column of $Y_{f}$ with itself. Consequently, at zeroth order $H_{f}$ has a single nonzero eigenvalue $y_{f_{3}}^{2}\sim 1$ (the $b$- or $t$-quark mass squared), while the lighter eigenvalues vanish.
The second eigenvalue $y_{f_{2}}^{2}$ emerges only after the rank-one piece is subtracted, exposing the $k=2$ contribution:
$y_{f_{2}}^{2}\sim\epsilon^{2Q(f^{c}_{2})}\times\left|c^{f}_{22}-\frac{c^{f}_{23}\,c^{f}_{32}}{c^{f}_{33}}\right|^{2},$ (135)
which reproduces the seesaw formula of Eqs. (120) and (128). The lightest eigenvalue $y_{f_{1}}^{2}$ requires a second such cancellation. The sequential bidiagonalization of the preceding subsections performs exactly these cancellations, one rotation at a time; the Hermitian-square approach makes them visible as the successive peeling of rank-one layers from $H_{f}$.
The advantage of the Hermitian-square viewpoint is that it exposes the exponent structure of $H_{f}$ as symmetric in the left-handed charges: $(H_{f})_{ij}\sim\epsilon^{Q(Q_{i})+Q(Q_{j})}$, independent of which column dominates the sum. This is why the left-handed mixing angles depend only on the left-handed charge differences $|Q(Q_{i})-Q(Q_{j})|$, as derived by both methods. The right-handed charges enter only through the eigenvalue hierarchy, appearing as the suppression $\epsilon^{2Q(f^{c}_{k})}$ of the successive rank-one layers. The two approaches are therefore complementary: the bidiagonalization is more efficient for extracting results, while the Hermitian-square construction reveals why the left–right factorization of the exponents translates into a clean separation of mixing angles (left charges) from mass eigenvalues (both charges).
C.4 CKM mixing angles
The CKM matrix $V=U_{uL}^{\dagger}U_{dL}$ receives contributions from both sectors. Since the shared left-handed doublet charges $Q(Q_{i})=(3,2,0)$ enter both $p^{u}$ and $p^{d}$ identically, the left-handed rotation angles have the same $\epsilon$-scaling in both sectors: $\theta_{ij}^{dL}\sim\theta_{ij}^{uL}\sim\epsilon^{|Q_{i}-Q_{j}|}$. The physical CKM elements therefore arise from interference between the two sectors:
$V_{ij}=(\theta_{ij}^{dL}-\theta_{ij}^{uL}\,e^{i\phi_{ij}})+\mathcal{O}(\theta^{2}),$ (136)
where $\phi_{ij}$ are relative phases from the $\mathcal{O}(1)$ coefficients.
The key observation is that although $\theta_{ij}^{dL}$ and $\theta_{ij}^{uL}$ share the same power of $\epsilon$, their $\mathcal{O}(1)$ prefactors differ because $Q(d^{c}_{j})\neq Q(u^{c}_{j})$. In the Fritzsch–Xing (FX) parameterization [7], the physical CKM elements are expressed in terms of the eigenvalue ratios:
$s_{d}$ $\equiv\sqrt{m_{d}/m_{s}}\sim\epsilon^{(37/9-7/3)/2}=\epsilon^{8/9},$ (137)
$s_{u}$ $\equiv\sqrt{m_{u}/m_{c}}\sim\epsilon^{(64/9-10/3)/2}=\epsilon^{17/9},$ (138)
$s$ $\equiv\sqrt{m_{c}m_{s}/(m_{t}m_{b})}\sim\epsilon^{(10/3+7/3)/2}=\epsilon^{17/6}.$ (139)
The Cabibbo angle then follows from the FX interference formula:
$|V_{us}|\simeq|s_{d}-s_{u}\,e^{i\phi_{\rm FX}}|\simeq s_{d}\sim\epsilon^{8/9}\simeq 0.225,$ (140)
since $s_{u}\sim\epsilon^{17/9}\ll s_{d}\sim\epsilon^{8/9}$. The remaining CKM elements scale as
$|V_{cb}|\sim\epsilon^{17/9}\simeq 0.042,\qquad|V_{ub}|\sim s_{d}\cdot s\sim\epsilon^{10/3}\simeq 0.004,$ (141)
in excellent agreement with experiment.
C.5 Relation to the Gatto–Sartori–Tonin formula
The classic Gatto–Sartori–Tonin (GST) relation [52] $\theta_{12}\simeq\sqrt{m_{d}/m_{s}}$, on which the FX parameterization relies, is strictly derived for symmetric or Hermitian mass matrices. The $B$-lattice textures are highly asymmetric: the left-handed charges $Q(Q_{i})=(3,2,0)$ differ markedly from the right-handed charges $Q(d^{c}_{j})=(\tfrac{10}{9},\tfrac{1}{3},0)$, so $Y_{d}\neq Y_{d}^{T}$. It is therefore important to clarify the status of the GST relation in this context.
In the perturbative diagonalization of Sec. VI, the 1122 left-handed mixing angle is governed purely by the left-handed charge difference:
$\theta_{12}^{dL}\sim\epsilon^{|Q(Q_{1})-Q(Q_{2})|}=\epsilon^{1}\simeq 0.187,$ (142)
independently of the right-handed charges. The GST mass-ratio expression, by contrast, involves both sets of charges through the eigenvalue ratio:
$\sqrt{m_{d}/m_{s}}=\epsilon^{(p_{d}-p_{s})/2}=\epsilon^{(Q_{1}-Q_{2}+Q^{d}_{c1}-Q^{d}_{c2})/2}=\epsilon^{8/9}\simeq 0.225.$ (143)
The two expressions are related by
$\theta_{12}^{dL}=\epsilon^{\,(Q_{1}-Q_{2}-Q^{d}_{c1}+Q^{d}_{c2})/2}\times\sqrt{m_{d}/m_{s}}=\epsilon^{1/9}\times\sqrt{m_{d}/m_{s}},$ (144)
where the correction factor $\epsilon^{1/9}\simeq 0.83$ is a direct measure of the left–right charge asymmetry. In the symmetric limit $Q(d^{c}_{j})\to Q(Q_{j})$ the exponent $(Q_{1}-Q_{2}-Q^{d}_{c1}+Q^{d}_{c2})/2$ vanishes, $\epsilon^{1/9}\to 1$, and the GST relation is recovered exactly.
The correction is numerically $\mathcal{O}(1)$ but lies on the ninths lattice, so it does not spoil the power-counting: both $\theta_{12}^{dL}\sim\epsilon$ and $\sqrt{m_{d}/m_{s}}\sim\epsilon^{8/9}$ are $\mathcal{O}(\epsilon)$ at leading order, and the distinction between the two is a subleading $\epsilon^{1/9}$ effect that is absorbed into the $\mathcal{O}(1)$ coefficients. The FX parameterization remains valid as an organizing scheme for the CKM magnitudes; the $\epsilon^{1/9}$ correction simply means that the $\mathcal{O}(1)$ prefactor relating $|V_{us}|$ to $\sqrt{m_{d}/m_{s}}$ is not unity but $\epsilon^{1/9}\simeq 0.83$ in the asymmetric $B$-lattice textures.
For the up sector the asymmetry is much larger: $Q(u^{c}_{1})-Q(u^{c}_{2})=25/9$ versus $Q(Q_{1})-Q(Q_{2})=1$, giving a correction exponent of 8/9-8/9 (opposite sign), so $\theta_{12}^{uL}\sim\epsilon\gg\sqrt{m_{u}/m_{c}}\sim\epsilon^{17/9}$. The GST relation is strongly violated in the up sector, reflecting the wide spacing of the up-type right-handed charges. This violation is harmless because the CKM is structurally down-dominated (Sec. VII.1): $|V_{us}|$ is controlled by $s_{d}\sim\epsilon^{8/9}$, not by $\theta_{12}^{uL}$.
Appendix D Illustrative Chain Coefficient Calculations
We illustrate the multi-messenger effective coefficient formula (23) with explicit numerical examples from the benchmark of Ref. [5]. In that construction, each sector has two sets of factorized shift matrices $\Delta^{f}_{ij}=\alpha^{f}_{i}+\beta^{f}_{j}$ and $\Delta^{\prime f}_{ij}=\alpha^{\prime f}_{i}+\beta^{\prime f}_{j}$, representing the additional suppression from the subleading messenger chains.
D.1 Down-sector $(2,3)$ entry
For the entry that controls $|V_{cb}|$, the benchmark shifts are $\alpha^{d}_{2}=0$, $\beta^{d}_{3}=4/9$, $\alpha^{\prime d}_{2}=1/9$, $\beta^{\prime d}_{3}=0$, giving
$\Delta^{d}_{23}=\tfrac{4}{9},\qquad\Delta^{\prime d}_{23}=\tfrac{1}{9}.$ (145)
With down-sector phases $\phi_{d}\simeq 2.86$, $\psi_{d}\simeq 3.54$:
$C^{d}_{23}$ $=1+e^{i\phi_{d}}\,\epsilon^{4/9}+e^{i\psi_{d}}\,\epsilon^{1/9}$
$=1+(-0.456+0.132i)+(-0.765-0.322i)$
$\simeq-0.221-0.190i.$ (146)
Thus $|C^{d}_{23}|\simeq 0.29$ and $\arg(C^{d}_{23})\simeq-139^{\circ}$. The destructive interference from the down-sector phases reduces the magnitude below unity but keeps it $\mathcal{O}(1)$; the nontrivial phase feeds into the CKM.
D.2 Up-sector $(2,3)$ entry
In the up sector the benchmark phases vanish ($\phi_{u}\simeq\psi_{u}\simeq 0$), and the shifts are $\Delta^{u}_{23}=5/9$, $\Delta^{\prime u}_{23}=1/9$, so the interference is purely constructive:
$C^{u}_{23}$ $=1+\epsilon^{5/9}+\epsilon^{1/9}$
$=1+0.394+0.830$
$\simeq 2.22.$ (147)
With all phases zero, $C^{u}_{23}$ is real and positive; $|C^{u}_{23}|\simeq 2.2$ is $\mathcal{O}(1)$. The absence of a CP-violating phase in the up sector is consistent with the “down-dominant” origin of the CKM phase.
D.3 Physical interpretation
The contrast between the two sectors is instructive. In the down sector, the phases $\phi_{d}$ and $\psi_{d}$ are close to $\pi$, causing partial cancellation and generating a large complex phase in $C^{d}_{ij}$. In the up sector, the benchmark phases vanish, yielding real coefficients of order 2. The CKM phase arises from the mismatch between these two patterns—the same adjacent-exponent interference mechanism of Sec. VII, now traced to its origin in the relative phases of the messenger chains.
Three aspects of this pattern warrant emphasis.
(i) The up-sector phases are inessential. As shown in Sec. VII.1, the CKM is structurally down-dominated: $s_{u}/s_{d}=\epsilon\simeq 0.19$, so the up-sector $\mathcal{O}(1)$ coefficients—including their phases—enter the CKM only as a $\sim\!19\%$ correction. Replacing $\phi_{u}=\psi_{u}=0$ with arbitrary $\mathcal{O}(1)$ phases changes $|V_{us}|$ by $\Delta|V_{us}|\lesssim s_{u}\simeq 0.04$ and shifts the Jarlskog invariant by a comparable fraction. The vanishing up-sector phases are adopted here for pedagogical clarity, not because the framework requires them.
(ii) Large CP violation is generic. The random-coefficient scan of Sec. VII.2 (Table 2) shows that $|\!\sin\delta|\gtrsim 0.3$ in $81\%$ of trials with fully random phases in both sectors. No tuning of phases—in either sector—is needed to obtain $\mathcal{O}(1)$ CP violation. The $\epsilon^{1/9}\simeq 0.83$ interference factor in Eq. (53) ensures that adjacent lattice exponents produce $\mathcal{O}(1)$ phase modulation for generic complex couplings.
(iii) One irreducible phase. The standard phase-counting theorem gives (n1)(n2)/2=1(n-1)(n-2)/2=1 physical CP phase for three generations. The multi-messenger phases $\phi_{f}$, $\psi_{f}$ are specific combinations of the UV Yukawa coupling phases that survive after all field rephasing; they do not represent additional parametric freedom. The framework’s structural claim is that the Jarlskog scaling $J\sim\epsilon^{19/3}\sin\delta$ is fixed by the lattice exponents, while the remaining free parameter $\sin\delta$ is generically $\mathcal{O}(1)$—it is not tuned to be either large or small.
The key point is that $|C^{f}_{ij}|$ remains $\mathcal{O}(1)$ in both cases: neither large hierarchies nor fine cancellations are required. The mass hierarchy is carried entirely by $\epsilon^{p^{f}_{ij}}$, while $C^{f}_{ij}$ provides the $\mathcal{O}(1)$ modulation and the physical CP phase.
D.4 Entry, propagation, and exit factors
The effective Yukawa amplitude for the $(i,j)$ entry factorizes into three physical pieces:
$(Y_{d})_{ij}=\underbrace{\lambda_{i}\,\epsilon^{Q(Q_{i})}}_{\text{entry}}\times\underbrace{\frac{\kappa_{1}\kappa_{2}\kappa_{3}}{M}\,\epsilon^{7/9}}_{\text{chain propagator}}\times\underbrace{\eta_{j}\,\epsilon^{Q(d^{c}_{j})}}_{\text{exit}}\times\,C^{d}_{ij}.$ (148)
The entry factor $\epsilon^{Q(Q_{i})}$ is the amplitude for left-handed quark generation $i$ to couple into the chain; the exit factor $\epsilon^{Q(d^{c}_{j})}$ is the amplitude for the chain to emit a right-handed quark of generation $j$. The chain propagator $(\kappa_{1}\kappa_{2}\kappa_{3}/M)\,\epsilon^{7/9}$ is common to all entries and is absorbed into the overall Yukawa normalization.
Example: $d$-quark Yukawa $(Y_{d})_{11}$. This entry connects the first-generation doublet to $d_{R}$ and controls the $d$-quark mass. The three factors are:
  • Entry: $\epsilon^{Q(Q_{1})}=\epsilon^{3}\simeq 6.5\times 10^{-3}$ (three flavon insertions at the left endpoint);
  • Chain: $(\kappa_{1}\kappa_{2}\kappa_{3}/M)\,\epsilon^{7/9}\simeq(\kappa_{1}\kappa_{2}\kappa_{3}/M)\times 0.27$ (hops $1/9+2/9+4/9$);
  • Exit: $\epsilon^{Q(d^{c}_{1})}=\epsilon^{10/9}\simeq 0.15$ (one flavon insertion plus a fractional hop at the right endpoint).
The charge-dependent suppression is $\epsilon^{Q(Q_{1})+Q(d^{c}_{1})}=\epsilon^{3+10/9}=\epsilon^{37/9}\simeq 10^{-3}$, which—multiplied by $|C^{d}_{11}|\sim\mathcal{O}(1)$—gives the observed ratio $m_{d}/m_{b}\simeq 10^{-3}$.
Example: $(Y_{d})_{23}$ entry (drives $V_{cb}$). Here:
  • Entry: $\epsilon^{Q(Q_{2})}=\epsilon^{2}\simeq 0.035$ (two flavon insertions);
  • Exit: $\epsilon^{Q(d^{c}_{3})}=\epsilon^{0}=1$ (no suppression—the $b_{R}$ couples directly to the chain end).
The charge-dependent part is $\epsilon^{2+0}=\epsilon^{2}\simeq 0.035$, and $|C^{d}_{23}|\simeq 0.29$ (from the destructive interference computed above), so $|(Y_{d})_{23}|\simeq 0.29\times\epsilon^{2}\simeq 0.010$. The unsuppressed exit factor for $b_{R}$ reflects the fact that the third-generation right-handed quark carries zero $Z_{9}$ charge and requires no flavon insertion to couple to the chain.
Acknowledgements.
VB gratefully acknowledges support from the William F. Vilas Estate.